### v2

# Vulnerability Webs:

Vulnerability Webs:
Systemic Risk in Software Networksthanks: The order of authors is alphabetical. fritzc@tcd.ie, co-pierre.georg@uct.ac.za, angelo.mele@jhu.edu, and mus47@psu.edu. This paper supersedes a previous version titled "A Strategic Model of Software Dependency Networks". We wish to thank Bryan Graham, Stephane Bonhomme, Louise Laage, Matt Jackson, Eric Auerbach for helpful suggestions. We also thank participants at 2024 Cowles Foundation Conference in Econometrics, ESIF 2024 Conference on ML+AI, Empirical Methods in Game Theory Workshop, MEG 2024, EC24, MGOW 2024, Network Science in Economics 2023, 2023 BSE Summer Forum Networks Workshop, SEA 2023, Econometrics of Peer Effects and Networks, and seminar participants at Georgetown University, University of Washington, Virginia Tech, Boise State University, and the University of Pennsylvania for helpful comments and suggestions. The authors acknowledge support from the German Research Foundation award DFG FR 4768/1-1 (CF), Ripple’s University Blockchain Research Initiative (CPG), the U.S. Department of Defense award ARO W911NF-21-1-0335 (MS, CF), and the U.S. National Science Foundation awards NSF DMS-1513644 and NSF DMS-1812119 (MS). Mele acknowledges support from an IDIES Seed Grant from the Institute for Data Intensive Engineering and Science at Johns Hopkins University and partial support from NSF grant SES 1951005.
Cornelius Fritz Trinity College Dublin – School of Computer Science and Statistics Co-Pierre Georg Frankfurt School of Finance and Management Angelo Mele Johns Hopkins University – Carey Business School Michael Schweinberger The Pennsylvania State University – Department of Statistics
Modern software development is a collaborative effort that re-uses existing code to reduce development and maintenance costs. This practice exposes software to vulnerabilities in the form of undetected bugs in direct and indirect dependencies, as demonstrated by the Crowdstrike and HeartBleed bugs. The economic costs resulting from such vulnerabilities can be staggering. We study a directed network of 52,897 software dependencies across 16,102 Python repositories, guided by a strategic model of network formation that incorporates both observable and unobservable heterogeneity. Using a scalable variational approximation of the conditional distribution of unobserved heterogeneity, we show that outsourcing code to other software packages by creating dependencies generates negative externalities. Modeling the propagation of risk in networks of software packages as an epidemiological process, we show that increasing protection of dependencies based on popular heuristics is ineffective at reducing systemic risk. By contrast, AI-assisted coding enables developers to replace dependencies with in-house code and reduces systemic risk.
Keywords: economics of software development, dependency graphs, strategic network formation, exponential random graphs
JEL Classification: D85, L17, C31, L86, O31
[…]
1 Introduction
On July 19, 2024 a faulty update to the software Crowdstrike—a security software adopted by many companies—created worldwide panic and disruptions, paralyzing domestic and international air traffic for hours, and disrupting operations at banks, hospitals, and hotel chains. [^1] See https://www.nytimes.com/2024/07/19/business/dealbook/tech-outage-crowdstrike-microsoft.html All told, the update affected 8.5 million Windows machines worldwide, each of which had to be manually rebooted to patch up the system and restart operations. The faulty update of the software Crowdstrike was not the first instance highlighting systemic risk in software systems, and it will not be the last: e.g., in 2014 the HeartBleed bug—a vulnerability of the OpenSSL authentication library used by many websites—compromised the data of millions of users of websites such as Eventbrite, OkCupid, in addition to disrupting operations at the FBI and hospitals. Such disruptions can generate staggering economic costs: Heartbleed, for example, resulted in at least \$currency-dollar\\$\$500 million in damages. According to the risk management company Parametrix, the disruption caused by Crowdstrike cost Fortune 500 companies \$currency-dollar\\$\$5.4 billion. [^2] See Parametrix (2024).
The risk of these disruptions is deeply embedded in the structure of software systems. Modern software development is a collaborative effort that produces sophisticated software, making extensive use of already existing software packages and libraries (Schueller et al., 2022). This results in a complex network of dependencies among software packages, best described as dependency graphs. [^3] Dependency graphs have been studied for many different programming languages using information from package managers. For an empirical comparison of the most common ones, see Decan et al. (2019). While this re-use of code produces significant efficiency gains for software developers (Lewis et al., 1991; Barros-Justo et al., 2018), it also increases the risk that a vulnerability in one software package renders a large number of other packages equally vulnerable. In line with this view, the number of vulnerabilities listed in the Common Vulnerabilities and Exposure Database has increased almost twelve-fold between 2001 and 2019.444The Common Vulnerabilities and Exposure database is a public reference for known information-security vulnerabilities which feeds into the U.S. National Vulnerability Database. See https://cve.mitre.org/. The reported number of incidents has increased from 3,591 for the three years from 1999 to 2001 to 43,444 for the three-year period from 2017 to 2019. A recent estimate by Krasner (2018) put the cost of losses from software failures at over \$1 trillion, up from \$59 billion in 2002 estimated by the National Institute of Standards and Technology (2002). [^5] Industry group Cybersecurity Ventures estimates that the damages incurred by all forms of cyber crime, including the cost of recovery and remediation, totalled USD 6 Trillion in 2021, and could reach USD 10.5 Trillion annually by 2025 (Source).
Because vulnerabilities can potentially spread from one software package to all packages depending on it, coders can create an externality when deciding to re-use code from existing packages rather than implementing the required functionality themselves. Whether this is the case and how large the externality is, is ultimately an empirical question.
We model coders’ decisions to create a dependency to another software package as an equilibrium game of network formation with observed and unobserved heterogeneity. Developers weigh costs and benefits of forming links and revise these links at random times, when there is a need for an update. Our equilibrium characterization provides the likelihood of observing a particular architecture of the software dependency network in the long run. Using data on open source software projects contained in Python’s software package manager Pypi, we obtain $52,897$ dependencies among $16,102$ repositories. We find evidence that a maintainer’s decision to allow a dependency exerts a negative externality on other maintainers. In principle, the presence of this externality discourages the formation of dependencies and partially contributes to the sparsity of the network. Our result implies that it is particularly important to ensure that such packages are free of vulnerabilities that potentially can affect a large number of other software packages.
We can use our model to study the effect of interventions that reduce vulnerabilities of libraries that are likely to affect a large subset of the software ecosystem. We assume that vulnerabilities, coding errors or bugs spread according to an epidemic susceptible-infected-recovered (SIR) model. Our simulations show that strategies to decrease systemic risk of contagion based on interventions that target nodes according to measures of centrality or expected risk of spread, are not very effective in mitigating the average systemic risk. On the other hand we observe that the introduction of AI-assisted coding may decrease the incentives to form dependencies through a decrease in costs of producing code. This process of decreasing the density of the network will mechanically decrease the average systemic risk. We show that even moderate productivity gains from AI translate into a significant flattening of the risk profile for the software network.
Open source software is an ideal laboratory to study software dependencies. Since the actual source code of a software package can be inspected and modified by others, code reuse is not only possible, it is encouraged. Modern software packages are not developed in isolation, but within an ecosystem of interdependent packages. A package manager keeps track of code dependencies and ensures that these are automatically resolved so that when a user installs one package, all of the package’s dependencies are also installed and automatically satisfied. There are many programming languages that provide software using package managers. We restrict our analysis to repositories of software packages written in the popular programming language Python and managed by the Pypi package manager.666 PyPI stands for Python Package Index and is Python’s official third-party software repository. See here for the official documentation. Python is a modern programming language that is very popular in the software development community. We use data from libraries.io, which, for each software repository includes a full list of all dependencies.
We model the development of software as a process where coders reuse existing code, creating a network of dependencies among software packages—organized in repositories, which can include several closely related software packages—in the process. This is typical for today’s prevalent object-oriented software development paradigm. Our model describes a system of $N$ individual software repositories, characterized by observable and unobservable types. Each repository is managed by a group of coders who decide which dependencies to form. Coders obtain a net benefit of linking to another package which depends on how active this dependency is maintained, how mature, popular, and large it is. We also allow a coders’ utility to be affected by a local, i.e. type-specific, externality: since dependent packages have dependencies themselves, they are susceptible to vulnerabilities imported from other packages. So we assume that coders care about the direct dependencies of the packages they link to. This specification excludes externalities that are more than two links away as in the network formation models of DePaula et al. (2018), Mele (2017) and Mele and Zhu (2023). The network of dependencies forms over time and in each period a randomly selected package needs an update, so the coders decide whether to form a link or update the software in-house. Before updating the link, the package receives a random match quality shock. We characterize the equilibrium distribution over networks as a mixture of exponential random graphs (Schweinberger and Handcock, 2015; Mele, 2022), which can be decomposed into within- and between-types contribution to the likelihood.
Model estimation is challenging because the likelihood depends on a normalizing constant that is infeasible to compute in large networks. Moreover, the model’s unobserved heterogeneity has to be integrated out in the likelihood, making direct computations infeasible even for moderately sized networks. To alleviate these problems, we resort to a novel approximate two-step approach. In the first step, we estimate the discrete unobservable types of the nodes by approximating their conditional distribution via a fast variational mean-field approximation first proposed for stochastic blockmodels (Vu et al., 2013). In the second step, we estimate the structural payoff parameters using a fast Maximum Pseudo-Likelihood Estimator (MPLE), conditioning on the estimated types (Babkin et al., 2020; Dahbura et al., 2021; Martínez Dahbura et al., 2023). We note that these methods were first developed for undirected graphs, and their application to directed networks necessitates non-trivial extensions of the computational machinery, providing an additional contribution of our work. [^7] Our methods are implemented in the scalable, open-source, and platform-independent R package bigergm, which is available on CRAN servers (R Core Team, 2024). Implementation details can be found in the Appendix.
Our main result is that coders exert a negative externality on other coders when creating a dependency on another package. Other coders are then also less likely to create a dependency with that package. As a result, the network is relatively sparse. The interdependence among libraries implies that there exists a risk that a vulnerability in a single package has large adverse consequences for the entire ecosystem. One example of how a vulnerability in one package impacted a significant part of critical web infrastructure is Heartbleed, the infamous bug in the widely used SSL/TLS cryptography library, resulting from an improper input variable validation (Durumeric et al., 2014). When Heartbleed was disclosed in 2014, it rendered up to 55% of secure webservers vulnerable to data theft, including of the Canadian Revenue Agency, and Community Health Systems, a large U.S. hospital chain. Our model provides a further argument for ensuring the security of highly interdependent software packages. Not only can a vulnerability in such a package affect a relatively large fraction of the entire ecosystem, because of the externality we identify, it is also likely that this fraction increases as time goes on.
Another important question in the study of network formation processes is whether linked packages are similar or dissimilar with respect to their covariates. We find evidence of similarity among the same type of software packages in terms of their Size, Popularityand Maturityas well as between different-type software packages in terms of their Popularity. The effect of similarity decreases with the (log of) the number of packages of the same unobservable type. We further find evidence of dis-similarity between different-type software packages in terms of their Maturity  and Size. In other words, mature, popular and large software packages of one type are likely to depend on similar packages of the same type, but on less mature and smaller software packages of another type. This is intuitive, because mature, popular, and large packages are likely to have a large user base with high expectations of the software and coders try to satisfy this demand by providing sophisticated functionality with the help of mature, popular, and large software packages of the same type. These packages are often built using a lot of low-level functionality from large but fairly generic libraries, which is why large popular and mature software packages of one type depend on larger, but less popular and less mature software packages of another type. One example of this is the inclusion of a payment gateway in an e-commerce application. The e-commerce application itself can be sophisticated and complex, like ebay is in the provision of their auction mechanism. For such an application it is particularly valuable to provide users with the ability to use a variety of different payment methods, including credit card, paypal, or buy now pay later solutions like Klarna. Our result also aligns with the recent trend in software development away from large monolithic applications and towards interconnected microservices (Traore et al., 2022).
Lastly, we adapt the susceptible-infected-recovered (SIR) model from epidemiology to examine the spread of vulnerabilities in the Python Pypi dependency graph. We measure a repository’s $k$-step systemicness as the number of downstream packages that are potentially rendered vulnerable by a vulnerability or bug in the upstream repository.
We then study the efficacy of targeted interventions to mitigate vulnerability spread, drawing parallels with vaccination strategies in public health. By focusing on securing the most critical nodes—determined according to in-degree, expected fatality or a combination of betweenness centrality and expected fatality—we assess the potential to significantly reduce the risk of vulnerability contagion.
Our findings suggest that protecting even ten percent of the most critical repositories may reduce the average systemic risk, but cannot completely flatten out the risk profile of the ecosystem. On the other hand, we argue that the introduction of AI-assisted coding may have a larger effect on reducing systemic risk. Recent work has shown that AI tools such as GitHub Copilot can increase productivity of software developers (Peng et al., 2023; Edelman et al., 2023; Noy and Zhang, 2023). In our model, this corresponds to a decrease of the cost of producing software in-house, without relying on creating dependencies or creating fewer dependencies. This can be modeled as an increase of the relative cost of a dependency. By simulating different magnitudes of the productivity improvement from AI-assisted coding, we show that the average systemic risk profile is flattened out by an increase of cost of merely five percent. The mechanism through which this risk reduction is achieved is the brute-force decrease in density for the network implied by increases in costs of dependencies. Overall, our results help better understand the various driving forces and motifs in software development and how they shape the network of software dependencies.
Our paper relates to several strands of literature in both economics and computer science. First, our paper contributes to a growing literature on open source software, which has been an interest of economic research. [^8] For an overview of the broad literature in the emerging field of digital economics, see Goldfarb and Tucker (2019). In an early contribution, Lerner and Tirole (2002) argue that coders spend time developing open source software—for which they are typically not compensated—as a signal of their skills for future employers. Likewise, one reason why companies contribute to open source software is to be able to sell complementary services. Open source projects can be large and complex, as Zheng et al. (2008) point out. They study dependencies in the Gentoo Linux distribution, and show that the resulting network is sparse, has a large clustering coefficient, and a fat tail. They argue that existing models of network growth do not capture the Gentoo dependency graph well and propose a preferential attachment model as alternative. [^9] In earlier work, LaBelle and Wallingford (2004) study the dependency graph of the Debian Linux distribution and show that it shares features with small-world and scale-free networks. However, LaBelle and Wallingford (2004) do not strictly check how closely the dependency graph conforms with either network growth model.
The package manager model is nowadays adopted by most programming languages which makes it feasible to use dependency graphs to study a wide variety of settings. Kikas et al. (2017), for example, study the structure and evolution of the dependency graph of JavaScript, Ruby, and Rust. The authors emphasize that dependency graphs of all three programming languages become increasingly vulnerable to the removal of a single software package. An active literature studies the network structure of dependency graphs (Decan et al., 2019) to assess the vulnerability of a software ecosystem (see, for example, Zimmermann et al., 2019).
These papers show the breadth of literature studying open source ecosystems and dependency graphs. However, the literature considers the network either as stochastic or even as static and given. In contrast, we model the formation of dependencies as coders’ strategic choice in the presence of various and competing mechanisms that either increase or reduce utility. [^10] Blume et al. (2013) study how possibly contagious links affect network formation in a general setting. While we do not study the consequences of the externality we identify for contagion, this is a most worthwhile avenue for future research. Perhaps the closest work is Boysel (2023), that uses data from a sample of repositories to estimate a structural model of network formation and maintainers effort. In his model, maintainers sequentially update their dependencies, then play a network game where their efforts best responses are a function of other maintainers efforts. While related, our approaches are different, since Boysel (2023) relies on sample panel data and the analysis focuses on the dynamics of myopic best-response. In our paper we focus on the long-run equilibrium and the cross-section of the dependencies among all repositories at a particular point in time.
The theoretical literature on strategic network formation has pointed out the role of externalities and heterogeneity in shaping the equilibrium networks (Bramoullé et al., 2016; Galeotti et al., 2006; Jackson and Wolinsky, 1996). Estimating strategic models of network formation is a challenging econometric task, because the presence of externalities implies strong correlations among links and multiple equilibria (Mele, 2017; Snijders, 2002; DePaula et al., 2018; Chandrasekhar, 2016; DePaula, 2017; Boucher and Mourifie, 2017; Graham, 2017, 2020). In this paper, we model network formation as a sequential process and focus on the long-run stationary equilibrium of the model (Mele, 2017, 2022). Because the sequential network formation works as an equilibrium selection mechanism, we are able to alleviate the problems arising from multiple equilibria.
Adding unobserved heterogeneity further complicates identification, estimation and inference (Schweinberger and Handcock, 2015; Graham, 2017; Mele, 2022). Other works have considered conditionally independent links without externalities (Graham, 2017, 2020; DePaula, 2017; Chandrasekhar, 2016), providing a framework for estimation and identification in random and fixed effects approaches. On the other hand, because link externalities are an important feature in this context, we move away from conditionally independent links, and model nodes’ unobserved heterogeneity as discrete types, whose realization is independent of observable characteristics and network, in a random effect approach. We can thus adapt methods of community discovery for random graphs to estimate the types, extending the work of Babkin et al. (2020) and in Dahbura et al. (2021); Martínez Dahbura et al. (2023) to accommodate for directed networks. Our two-steps method scales well to large networks, thus improving the computational challenges arising in estimation of these complex models (Boucher and Mourifie, 2017; Vu et al., 2013; Bonhomme et al., 2019).
Our model is able to estimate the magnitude of externalities as well as homophily (Currarini et al., 2010; Jackson, 2008; Chandrasekhar, 2016), the tendency of individuals to form links to similar individuals. Furthermore, our model is able to detect heterophily (or competition) among maintainers. More specifically, we allow the homophily to vary by unobservables, while in most models the homophily is estimated only for observable characteristics (Currarini et al., 2010; Schweinberger and Handcock, 2015; DePaula et al., 2018; Chandrasekhar, 2016; Graham, 2020).
2 A network description of code
2.1 The open source software paradigm
The goal of computer programs is to implement algorithms on a computer. An algorithm is a terminating sequence of operations which takes an input and computes an output using memory to record interim results. [^11] Memory to record interim instructions is sometimes called a “scratch pad”, in line with early definitions of algorithms which pre-date computers. See, for example, Chapter 1 of Knuth (1997). We use the term broadly to include algorithms that rely heavily on user inputs and are highly interactive (e.g. websites, spreadsheet and text processing software, servers). Algorithms are implemented on a computer using code. Formally:
Definition 1.
Code is a sequence of operations and arguments that implement an algorithm. A computer program is code that can be executed by a computer system.
In order to execute a program, a computer provides resources—processing power and memory—and resorts to a compiler or interpreter, which in themselves are computer programs. [^12] The term ”compiler” was coined by Hopper (1952) for her arithmetic language version 0 (A-0) system developed for the UNIVAC I computer. The A-0 system translated a program into machine code which are instructions that the computer can execute natively. Early compilers were usually written in machine code or assembly. Interpreters do not translate program code into machine code, but rather parse it and execute the instructions in the program code directly. Early interpreters were developed roughly at the same time as early compilers, but the first widespread interpreter was developed in 1958 by Steve Russell for the programming language LISP (see McCarthy (1996).
Software developers, which we refer to as coders, use programming languages to implement algorithms. Today, the dominant modern software development paradigm is object-oriented programming (OOP). [^13] For a principal discussion of object-oriented programming and some differences to procedural programming, see Abelson et al. (1996). Kay (1993) provides an excellent historical account of the development of early object-oriented programming. Under this paradigm, code is developed primarily in classes, which contain data in the form of variables and data structures as well as code in the form of procedures. [^14] Also called functions, or methods in the OOP paradigm. Classes can communicate with one another via procedures using calls and returns.
Definition 2.
A class is a code template defining variables, procedures, and data structures. An object is an instance of a class that exists in the memory of a computer system.
Classes can interact in two ways. First, in the traditional monolithic software architecture, widely used for enterprise software like the one developed by SAP, for operating systems like Microsoft’s Windows, and even in earlier versions of e-commerce platforms like Amazon, individual components cannot be executed independently. In contrast, many modern software projects use a microservices software architecture, which is a collection of cohesive, independent processes, interacting via messages. Both architectures result in software where individual pieces of code depend on other pieces, either within a single application or across various microservices. These dependencies form a network which we formalize in the next section.
2.2 Dependency Graphs
Most open source software is organized by package managers like PyPi for the Python programming language that standardize and automate software distribution. To make the installation of complex software projects easier for users, package managers keep track of a package’s dependencies and allow users to install these alongside the package they are installing. The existence of package managers and the level of automation they provide is necessary because modern software is frequently updated and different versions of the same package are not always compatible. Packages are logically grouped into repositories, controlled and managed by a maintainer. [^15] Maintainers can also be organizations. Some repositories are also controlled by a group of maintainers, but for our purposes, this is immaterial.
[…]
There are two reasons for constructing the dependency graph on the repository rather than the package level. First, information about the popularity of a piece of software is relevant for maintainers linking decisions, but only created on the repository level. [^17] For example, users can “star” a repository on GitHub if they find it particularly useful. And second, repositories provide a logical grouping of closely related software packages. Alternatively, we could construct the dependency graph on the level of individual software packages. However, the breadth of code and functionality between packages can be much smaller than the breadth of code and functionality within a package. [^18] Yet another alternative is to study call graphs arising from procedures within the same software package (see, for example, Grove et al., 1997). But these are state-dependent, i.e., dependencies arise during runtime and depending on how the package is executed and interacted with. Conveniently, the website libraries.io provides information collected from various open source software package managers, including PyPi. [^19] The website libraries.io obtains this data by scraping publicly available repositories hosted on GitHub, GitLab, and Bitbucket. The data provided includes information on projects–essentially the same as a PyPi package–as well as repositories and for each repository a list of all dependencies, defined on the project level. Since each project belongs to exactly one repository, we construct dependencies among repositories from the data provided.
On the repository level, this data includes information about the size of the repository in kB (Size), its popularity, measured as the number of stars it has on the website hosting the repository (Popularity), and as an alternative measure of popularity the number of contributors, i.e. the number of individual coders who have added code to a repository, raised or answered an issue, or reviewed code submitted by others. [^20] We restrict ourselves to repositories that have a size above the 5% percentile of the size distribution, and that have more than $2$ stars, but no more stars than the 95% percentile of the distribution of stars.
#Nodes #Edges #Components #Nodes (LCC) #Edges (LCC)
$17,095$ $53,498$ $422$ $16,102$ $52,897$
Table 1: Network statistics for the dependency graph of the PyPi ecosystem.
Mean Std.Dev. Min. p5 Median P95 Max.
In-Degree $3.29$ $50.49$ 00 00 00 $6$ $4,210$
Out-Degree $3.29$ $5.0$ 00 00 $2$ $12$ $95$
Table 2: Distribution of in- and out-degree for the largest weakly connected component of the repo-based dependency graph of the PyPi ecosystem with $N=16,102$ nodes.
Figure 1: Dependency graph of the PyPi ecosystem.
Table LABEL:Table::NetworkStats provides a high-level overview of the repository-based dependency graph constructed in this way. While there are $91$ weakly connected components, the largest weakly connected component covers almost all (over $99\%$) nodes and edges. This means that we still cover all relevant dynamics even by only focusing on the largest weakly connected component of the dependency graph.
Furthermore, Table 2 shows that there is a lot of heterogeneity in how interconnected repositories are. The standard deviation of the out-degree, i.e. how many dependencies a given repository uses, is almost twice as large as the mean. Therefore, we observe substantial heterogeneity among repositories, with the median repository using $2$ dependencies and the maximum being $95$ dependencies. This heterogeneity is even more pronounced for the in-degree, i.e. how often a given repository is a dependency for another repository. The standard deviation is more than ten times as large as the mean; the median is 00, while the maximum is $14,585$.
In Table LABEL:Tab::TopRepos we show the $10$ most depended-on packages (i.e. with the highest in-degree) The repository with highest in-degree is six, which facilitates compatibility between python version 2 and 3 (where some major updates have led to serious incompatibility issues). The second-highest in-degree repository is pytest-cov, which is used to check a code’s test coverage, which is a crucial step in modern software development. Likewise, the third-most commonly linked software package is mock, which is also used for software testing.
Covariates
Repo name In-Degree Popularity Size Maturity
benjaminp/six 4,210 579 1,749 1,106
pytest-dev/pytest-cov 2,584 643 777 2,097
testing-cabal/mock 2,002 421 1,306 1,728
pypa/wheel 1,572 156 1,302 902
kjd/idna 1,336 106 507 2,422
pytest-dev/pluggy 952 364 471 1,721
pypa/packaging 782 141 540 2,069
pallets/markupsafe 775 282 175 3,492
PyCQA/mccabe 767 242 82 2,518
coveralls-clients/coveralls-python 761 406 328 2,532
Table 3: List of the ten most depended-on repositories for the PyPi ecosystem.
The largest weakly connected component of the repo-based dependency graph is shown in Figure LABEL:Fig::RustDependencyGraph, where we apply the community detection algorithm of Blondel et al. (2008) to color-code packages in the same community. The three largest communities contain about $10-11\%$ of all repositories each.
2.3 Observables: Covariates
In addition to the software dependency data, we obtain additional data from the from libraries.io website for each package (e.g. https://libraries.io/pypi/pandas), which we show for the ten most-depended on repositories in Table LABEL:Tab::TopRepos. We use three types of covariates, summarized in a vector of observable attributes $\bm{x}_{i}$. First, we use a repository’s Size of all code in the repository, measured in Kilobytes. Second, Popularity is measured as the number of stars–an expression of how many people like a repository or find it useful–a repository received on the website hosting it, e.g., GitHub. And, third, the time in days between the first release of a repository and the reference date of 13 January 2020 is denoted Maturity.
Mean Std.dev Min p5 Median p95 Max
Size [kB] 7,729.8 52,557.84 6 23 273 27,627 3,484,077
Popularity 70.49 133.35 3 3 16 364 853
Maturity 1384.6 838.14 4 248 1250 2960 4270
Table 4: Distribution of covariates
To facilitate the estimation of our model, we create a categorical variable for each of our covariates using quartiles of the distribution. The reason for discretizing variables is computational, as our algorithm is based on stochastic block models with discrete types (Bickel et al., 2013; Vu et al., 2013). If the covariates are discrete, the estimation engine of stochastic block models can be adapted to estimate the unobserved heterogeneity (Vu et al., 2013; Babkin et al., 2020; Dahbura et al., 2021). Otherwise, when the covariates are continuous, the computational costs are prohibitive for such large networks. [^21] We provide more details in the appendix; see also Babkin et al. (2020) and Dahbura et al. (2021).
3 Model
We model software development as a dynamic process, assuming that the developers update their packages to add or improve their functionality. Maintainers can update their packages by either adding functionalities to their own packages or by creating dependencies to other packages that have the desired functionalities.
3.1 Setup
Each package $i$ and its developer team is characterized by a vector of $P$ observable attributes $\bm{x}_{i}\in\mathbb{R}^{P}$, which are stored in a matrix $\bm{x}\in\mathbb{R}^{N\times P}$. We assume that each of the $N$ packages belongs to one of $K\geq 2$ types, unknown to researchers but known to other developers. The type of a software package may be its purpose, which may be known to developers but may not be known to researchers, or may correspond to unobservable qualities of the code and its developer team. To indicate the type $k$ of package $i$, we define $z_{ik}\coloneqq 1$ and $z_{il}\coloneqq 0$ for all $l\neq k$ if package $i$ is of type $k$, and write $\bm{z}_{i}\coloneqq(z_{i1},\ldots,z_{iK})$. We assume that the unobservables are independent and identically distributed according to a multinomial distribution:
[…]
where $\bm{\eta}\coloneqq(\eta_{1},\ldots,\eta_{K})$ is a vector of $K$ probabilities summing to $1$. We represent the software dependency network $\bm{g}$ by the adjacency matrix, with elements $g_{ij}=1$ if package $i$ depends on package $j$ and $g_{ij}=0$ otherwise. In general, software dependency networks are directed: If package $i$ depends on package $j$, then package $j$ need not depend on package $i$.
The utility that developers of package $i$ receive from the software dependency network $\bm{g}$, with observables $\bm{x}$, unobservables $\bm{z}$, and parameters $\bm{\theta}\coloneqq(\bm{\alpha},\,\bm{\beta},\,\gamma)$, is of the form
$U_{i}(\bm{g};\,\bm{x},\,\bm{z},\,\bm{\theta})\;\coloneqq\;\sum_{j=1}^{N}g_{ij}u_{ij}(\bm{\alpha},\,\bm{\beta})+\sum_{j=1}^{N}\,\sum_{r\neq i,j}^{N}g_{ij}g_{jr}v_{ijr}(\gamma)+\sum_{j=1}^{N}\,\sum_{r\neq i,j}^{N}g_{ij}g_{ir}w_{ijr}(\gamma).$ (2)
The function $u_{ij}(\bm{\alpha},\,\bm{\beta})\coloneqq u(\bm{x}_{i},\,\bm{x}_{j},\bm{z}_{i},\,\bm{z}_{j},\bm{\alpha},\,\bm{\beta})$ is the direct utility of linking package $i$ to package $j$. It is a function of observables ($\bm{x}_{i},\,\bm{x}_{j}$), unobservables ($\bm{z}_{i},\,\bm{z}_{j}$), and parameters ($\bm{\alpha},\,\bm{\beta}$). This represents the benefit of a dependency of package $i$ on package $j$, net of costs: e.g., the developers will have to audit the code of package $j$ and determine its quality, and have to maintain the dependency on package $j$. The cost may includes modification and adaptation of the code, because the developers have to be able to use the functions and methods available in package $j$. Additional costs may arise from relying on other packages, because other packages may contain undetected bugs. Finally, relying on existing code is convenient but may decrease the incentives of developers to maintain and improve coding skills, which in turn can undermine the future productivity of developers and can hence be viewed as a cost.
Motivated by the application to software dependency networks, we parameterize $u_{ij}$ as follows:
$u(\bm{x}_{i},\bm{x}_{j},\,\bm{z}_{i},\bm{z}_{j},\,\bm{\alpha},\,\bm{\beta})\;\coloneqq\;\begin{cases}u_{w}(\bm{x}_{i},\bm{x}_{j},\,\bm{z}_{i},\bm{z}_{j},\,\bm{\alpha}_{w},\,\bm{\beta}_{w})\ &\text{ if both $i$ and $j$ are of type $k$}\vspace{.25cm}\\ u_{b}(\bm{x}_{i},\bm{x}_{j},\,\bm{z}_{i},\bm{z}_{j},\,\bm{\alpha}_{b},\,\bm{\beta}_{b})\ &\text{otherwise,}\end{cases}$ (3)
where the utility for dependencies within types $u_{w}(\bm{x}_{i},\bm{x}_{j},\,\bm{z}_{i},\bm{z}_{j},\,\bm{\alpha}_{w},\,\bm{\beta}_{w})$ is
$\begin{array}[]{lllll}u_{w}(\bm{x}_{i},\bm{x}_{j},\,\bm{z}_{i},\bm{z}_{j},\,\bm{\alpha}_{w},\,\bm{\beta}_{w})&\coloneqq&\alpha_{w1}+\alpha_{w1}^{\ell}\,\log(N_{k})+\displaystyle\sum_{p=1}^{P}\beta_{wp}\,h_{ijp}(\bm{x}_{i},\bm{x}_{j})\\ &+&\displaystyle\sum_{p=1}^{P}\beta_{wp}^{\ell}\,h_{ijp}(\bm{x}_{i},\bm{x}_{j})\,\log(N_{k}),\end{array}$ (4)
and the utility for dependencies between types $u_{b}(\bm{x}_{i},\bm{x}_{j},\,\bm{z}_{i},\bm{z}_{j},\,\bm{\alpha}_{b},\,\bm{\beta}_{b})$ is
$\begin{array}[]{lllll}u_{b}(\bm{x}_{i},\bm{x}_{j},\,\bm{z}_{i},\bm{z}_{j},\,\bm{\alpha}_{b},\,\bm{\beta}_{b})&\coloneqq&\alpha_{b}+\displaystyle\sum_{p=1}^{P}\beta_{bp}\,h_{ijp}(\bm{x}_{i},\bm{x}_{j}).\end{array}$ (5)
The terms in the within- and between-type utilities can be interpreted as follows: The intercept $\alpha_{w1}\in\mathbb{R}$ represents the fixed cost of forming a dependency; the term $\alpha_{w2}\,\log(N_{k})$ with weight $\alpha_{w2}\in\mathbb{R}$ is the part of the cost that depends on the size $N_{k}$ of type $k$; [^22] In principle, the specification can be extended to incorporate aggregated terms at the type-level. In this sense, $\log(N_{k})$ is one among other examples of terms that can be used to make the model more flexible. For instance, one may include the average of observed characteristics of type $k$ as one of the covariates. We do not pursue this specification here, given the very sparse nature of the software dependency network. and $h_{ijp}(\bm{x}_{i},\bm{x}_{j})\coloneqq\bm{1}\{{x}_{ip}={x}_{jp}\}$ is an indicator function, which is equal to $1$ if the $p$th covariate of $i$ and $j$ matches and is 00 otherwise. Size-dependent terms of the form $\log(N)$ were suggested by Krivitsky et al. (2011), motivated by invariance considerations (encouraging the expected number of links to be invariant to $N$), and adapted to observed types by Krivitsky et al. (2023) and to unobserved types by Babkin et al. (2020). The micro-behavioral foundations of size-dependent terms of the form $\log(N)$ were studied by Butts (2019). The vectors $\bm{\alpha}\coloneqq(\alpha_{w1},\,\alpha_{w2},\,\alpha_{b})\in\mathbb{R}^{3}$ and $\bm{\beta}\coloneqq(\beta_{w1},\,\ldots,\,\beta_{wP},\,\beta_{b1},\,\ldots,\,\beta_{bP},\beta_{w1}^{\ell},\,\ldots,\,\beta_{wP}^{\ell})\in\mathbb{R}^{3P}$ are parameters to be estimated.
The second and third term of utility function (2) are externalities generated by linking to package $j$, because package $j$ may be linked to other packages $r$ as well. This means that the developers of package $i$ need to check the quality and features of the code in those packages $r$ as well, to ensure that those packages are free of bugs or vulnerabilities, and compatible with package $i$. On the other hand, any update of package $r$ may compromise compatibility with packages $i$ and $j$, increasing the cost of maintaining package $i$. In short, $w_{ijr}$ measures the net benefits of this externality when developers of package $i$ form a dependency to package $j$. In addition, if the developers of package $i$ create a dependency on package $j$, it may compromise compatibility with other packages $r$ that depend on package $i$. This is accounted for by the third term $v_{ijr}$ in utility function (2). We assume that the second and third term in the utility are of the form
$v_{ijr}(\gamma)\;=\;w_{ijr}(\gamma)\;\coloneqq\;\begin{cases}\gamma&\text{ if $i$, $j$, $r$ are all of type $k$}\\ 0&\text{otherwise,}\end{cases}$ (6)
where $\gamma\in\mathbb{R}$.
The choice of the utility functions $U_{i}$ guarantees that all externalities are local, in the sense that all externalities are limited to packages of the same type. Local externalities make sense in large networks, because developers are unlikely to know the functionalities of thousands of other packages and, more likely than not, will create dependencies to other packages of the same type, which developers are most familiar with. In addition, local externalities have probabilistic, statistical, and computational advantages. First, local externalities ensure desirable properties of models for large networks (Schweinberger and Handcock, 2015; Schweinberger et al., 2020). Second, local externalities facilitate theoretical guarantees for statistical procedures (Schweinberger and Stewart, 2020; Schweinberger, 2020; Stewart and Schweinberger, 2023; Stewart, 2024). Third, local externalities enable local computing on subnetworks of large networks, which facilitates parallel computing on a massive scale (Babkin et al., 2020).
3.2 Equilibrium
Software development is a dynamic process. We assume that time is discrete and that at each time point $t=1,2,\ldots$ a single package is updated. In other words, maintainers do not update software packages continuously, but certain events – unobservable to researchers – trigger the need for an update. For example, if package $i$ depends on package $j$, then an update of package $j$ may require an update of package $i$, or a request by users of package $i$ may trigger an update of package $i$. The update can be produced in-house or by forming a dependency to some other package $j$ containing the same functionality. Since researchers may not be able to observe the timing of such updates, we model updates as a discrete-time Markov process (Norris, 1997).
The first ingredient if the discrete-time Markov process is the update process that determines which package $i$ requires an update and which dependency $j$ is considered to implement the update. We assume that the probability that package $i$ requires an update and the developers of package $i$ consider creating, maintaining, or removing a dependence to package $j$ is $\rho(\bm{g}^{t-1},\bm{x}_{i},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j})$, where $\bm{g}^{t-1}$ is the network at time $t-1$. So the update process may depend on the existing network $\bm{g}^{t-1}$ as well as observables ($\bm{x}_{i}$, $\bm{x}_{j}$) and unobservables ($\bm{z}_{i}$, $\bm{z}_{j}$). A simple example is an update process that selects an ordered pair of packages $i$ and $j$ at random. In other words, a package $i$ is selected with probability $1/n$ and requires an update. Conditional on the event that package $i$ requires an update, the developers of package $i$ consider updating the possible dependency of package $i$ on package $j$ with probability $1/(n-1)$. As a result, the probability that package $i$ requires an update and the developers of package $i$ consider updating a (possible) dependency of package $i$ on package $j$ is
$\rho\left(\bm{g}^{t-1},\bm{x}_{i},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j}\right)\;\;=\;\;\dfrac{1}{n(n-1)}.$
A second example is an update process that allows the conditional probability that an update of the (possible) dependency of package $i$ on package $j$ is proposed to depend on the distance $d(x_{i},x_{j})$ between packages $i$ and $j$:
$\rho\left(\bm{g}^{t-1},\bm{x}_{i},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j}\right)\;\;\propto\;\;\frac{1}{n}\exp\left(-d(\bm{x}_{i},\bm{x}_{j})\right),$
where $d(\bm{x}_{i},\bm{x}_{j})$ is a function of observables ($\bm{x}_{i}$, $\bm{x}_{j}$) satisfying reflexity, symmetry, and the triangle inequality. In other words, $d(\bm{x}_{i},\bm{x}_{j})$ is a distance function that quantifies the distance between packages $i$ and $j$ based on observable characteristics $\bm{x}_{i}$ and $\bm{x}_{j}$ of packages $i$ and $j$.
[…]
$\rho\Big{(}\bm{g}^{t-1},\bm{x}_{i},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j}\Big{)}\;\;=\;\;\rho\Big{(}\bm{g}_{-(i,j)}^{t-1},\bm{x}_{i},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j}\Big{)}.$
Second, the update probabilities $\rho(\bm{g}_{-(i,j)}^{t-1},\bm{x}_{i},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j})$ are strictly positive for all ordered pairs of packages $(i,j)$:
[…]
This implies that each possible dependency can be proposed with a positive probability $\rho$, no matter how small. These assumptions help ensure that in the long-run equilibrium these probabilities do not affect the distribution of networks (Mele, 2017).
While the update process is stochastic, the decision to create, maintain, or remove a possible dependency of package $i$ on package $j$ is a strategic decision. Conditional on the opportunity to update the dependency of package $i$ on package $j$, the developers of package $i$ make a decision about whether to create, maintain, or remove a dependency on package $j$ based on utility function $U_{i}(\bm{g};\,\bm{x},\bm{z},\bm{\theta})$. Before the developers of package $i$ make a decision about the (possible) dependency on package $j$, the developers receive random shocks $\varepsilon_{ij0}\in\mathbb{R}$ and $\varepsilon_{ij1}\in\mathbb{R}$ to the perceived utility of a dependency of package $i$ on package $j$ ($g_{ij}^{t}=1$) and the perceived utility of writing code in-house ($g_{ij}^{t}=0$). After receiving random shocks $\varepsilon_{ij0}\in\mathbb{R}$ and $\varepsilon_{ij1}\in\mathbb{R}$, the developers of package $i$ create a dependency on package $j$ if the perceived utility of the dependency of package $i$ on package $j$ ($g_{ij}^{t}=1$) exceeds the utility of writing code in-house ($g_{ij}^{t}=0$):
$U_{i}\left(g_{ij}^{t}=1,\,\bm{g}_{-(i,j)}^{t-1};\,\bm{x},\bm{z},\bm{\theta}\right)+\varepsilon_{ij1}^{t}\;\geq\;U_{i}\left(g_{ij}^{t}=0,\,\bm{g}_{-(i,j)}^{t-1};\,\bm{x},\bm{z},\bm{\theta}\right)+\varepsilon_{ij0}^{t},$ (7)
where $\bm{g}_{-(i,j)}^{t-1}$ refers to network $\bm{g}^{t-1}$ excluding the possible dependency $g_{ij}$ of package $i$ on package $j$. The random shocks to payoffs $\varepsilon_{ij0}^{t}\in\mathbb{R}$ and $\varepsilon_{ij1}^{t}\in\mathbb{R}$ capture the fact that unexpected events can occur when developing code in-house ($\varepsilon_{ij0}$) and when creating a dependency ($\varepsilon_{ij1}$), e.g., changes in the composition of the developer team or additional information about package $j$. We assume that the random shocks to payoffs $\varepsilon_{ij0}$ and $\varepsilon_{ij1}$ do not depend on time, are independent across pairs of packages $(i,j)$, and follow a logistic distribution, which is a standard assumption in the literature on discrete choice models (Mele, 2017; Graham and dePaula, 2020; Boucher and Mourifie, 2017). As a consequence, when the opportunity to update the dependency of package $i$ on package $j$ arises, the conditional choice probability of a dependency of package $i$ on package $j$ is
$\mathbb{P}\left(g_{ij}^{t}=1\ |\ \bm{g}_{-(i,j)}^{t-1};\,\bm{x},\bm{z},\bm{\theta}\right)=\frac{\exp\left(U_{i}(g_{ij}^{t}=1,\,\bm{g}_{-(i,j)}^{t-1};\,\bm{x},\bm{z},\bm{\theta})-U_{i}(g_{ij}^{t}=0,\,\bm{g}_{-(i,j)}^{t-1};\,\bm{x},\bm{z},\bm{\theta})\right)}{1+\exp\left(U_{i}(g_{ij}^{t}=1,\,\bm{g}_{-(i,j)}^{t-1};\,\bm{x},\bm{z},\bm{\theta})-U_{i}(g_{ij}^{t}=0,\,\bm{g}_{-(i,j)}^{t-1};\,\bm{x},\bm{z},\bm{\theta})\right)}.$ (8)
As shown in Mele (2022) (with unobservables) and Butts (2009) and Mele (2017) (without unobservables), the sequence of networks generated by the process described above is a discrete-time Markov process (Norris, 1997), conditional on the unobservable types $\bm{z}$.
The discrete-time Markov chain described above is irreducible and aperiodic and hence admits a unique limiting distribution $\pi$, which is the stationary distribution of the Markov chain (Norris, 1997). The stationary distribution $\pi$ can be represented in closed form as an exponential-family distribution (Mele, 2017) and is given by
$\pi(\bm{g};\,\bm{x},\,\bm{z},\,\bm{\theta})$$\coloneqq$ $\left(\prod_{k=1}^{K}\frac{e^{Q_{kk}(\bm{g}_{kk};\,\bm{x},\,\bm{z},\,\bm{\theta})}}{c_{kk}(\bm{x},\,\bm{z};\,\bm{\theta})}\right)\;\left(\prod_{l\neq k}^{K}\,\prod_{i=1}^{N_{k}}\,\prod_{j=1}^{N_{l}}\frac{e^{u_{ij}(\bm{\alpha}_{b},\,\bm{\beta}_{b})}}{1+e^{u_{ij}(\bm{\alpha}_{b},\,\bm{\beta}_{b})}}\right).$ (9)
The potential function $Q_{kk}$ of the within-type-$k$ distribution is defined as
$\begin{array}[]{lllll}Q_{kk}\left(\bm{g}_{kk};\,\bm{x},\,\bm{z},\,\bm{\theta}\right)&\coloneqq&\displaystyle\sum_{i=1}^{N}\,\sum_{j\neq i}^{N}u_{ij}\left(\bm{\alpha}_{w},\bm{\beta}_{w},\bm{x_{i}},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j}\right)g_{ij}z_{ik}z_{jk}\vspace{.25cm}\\ &+&\gamma\,\displaystyle\sum_{i=1}^{N}\,\sum_{j\neq i}^{N}\,\sum_{r\neq i,j}^{N}g_{ij}g_{jr}z_{ik}z_{jk}z_{rk},\end{array}$ (10)
while the normalizing constant $c_{kk}(\bm{x},\,\bm{z};\,\bm{\theta})$ of the within-type-$k$ distribution is defined as
$c_{kk}(\bm{x},\,\bm{z},\,\bm{\theta})$$\coloneqq$ $\displaystyle\sum_{\bm{\omega}\,\in\,\mathcal{G}_{kk}}e^{Q_{kk}(\bm{\omega};\;\bm{x},\,\bm{z},\,\bm{\theta})},$ (11)
where the sum is over all possible networks $g_{kk}^{\prime}\in\mathcal{G}_{kk}$ of type $k$.
The limiting distribution $\pi$ of the discrete-time Markov chain represents the long-run distribution of the software development process, conditional on the types $\bm{z}$ of packages. The first term in (9) represents the likelihood of dependencies among packages of the same type, while the second term represents the likelihood of dependencies between packages of different types.
The stationary distribution $\pi$ implies that the incentives of developers are captured by a potential function of the form
$Q(\bm{g},\bm{x},\bm{z},\bm{\theta})$ $=$ $\sum_{k=1}^{K}Q_{kk}\left(\bm{g}_{kk};\,\bm{x},\,\bm{z},\,\bm{\theta}\right)+\sum_{k=1}^{K}\sum_{l\neq k}Q_{kl}\left(\bm{g}_{kl};\,\bm{x},\,\bm{z},\,\bm{\theta}\right),$ (12)
where $Q_{kk}\left(\bm{g}_{kk};\,\bm{x},\,\bm{z},\,\bm{\theta}\right)$ is defined in (10) while $Q_{kl}\left(\bm{g}_{kl};\,\bm{x},\,\bm{z},\,\bm{\theta}\right)$ is defined by
$Q_{kl}\left(\bm{g}_{kl};\,\bm{x},\,\bm{z},\,\bm{\theta}\right)\;\;\coloneqq\;\;\sum_{i=1}^{N}\sum_{j=1}^{N}g_{ij}\,u_{b}(\bm{x}_{i},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j},\bm{\alpha_{b}},\bm{\beta}_{b})\,z_{ik}\,z_{jl}.$ (13)
In this software development process, each developer team optimally responds to the existing software dependency network. The Nash equilibria of the underlying potential game of the software development process are the networks that maximize the potential function $Q$ (Mele, 2017, 2022; Monderer and Shapley, 1996; Chandrasekhar, 2016; Blume et al., 2013). These are networks such that no developer team is willing to form an additional dependency or delete an existing dependency, when the need arises to update package dependencies. Upon inspecting the stationary distribution $\pi$, it is evident that the Nash equilibria have the highest probability of being observed, because the Nash equilibria maximize $Q$ and hence maximize $\pi$.
Having said that, Nash networks will be inefficient in the presence of externalities. Consider the welfare function defined by the sum of the individual maintainers utilities
$\begin{array}[]{llllll}W(\bm{g},\bm{x},\bm{z},\bm{\theta})&\coloneqq&\displaystyle\sum_{i=1}^{N}U_{i}(\bm{g},\bm{x},\bm{z},\bm{\theta})\vspace{.25cm}\\ &=&Q(\bm{g},\bm{x},\bm{z},\bm{\theta})\,+\,\gamma\,\displaystyle\sum_{k=1}^{K}\sum_{i=1}^{N}\,\sum_{j\neq i}^{N}\,\sum_{r\neq i,j}^{N}g_{ij}g_{ri}z_{ik}z_{jk}z_{rk},\end{array}$ (14)
which is not equal to the potential function $Q(\bm{g},\bm{x},\bm{z},\bm{\theta})$ when $\gamma\neq 0$, so the networks that maximize welfare may not maximize the potential function and are hence not Nash equilibria.
4 Estimation and Empirical Results
We perform the estimation under the assumption that the observed software dependency network is a sample from the stationary distribution $\pi$ of the discrete-time Markov process described in Section 3.2. This implies that we can estimate the structural parameters based on a single observation of the software dependency network. We sketch a scalable two-step estimation algorithm and refer the interested reader to the appendix for more technical details.
4.1 Estimation
We introduce a scalable approach to estimating directed network formation models with observable and unobservable types. While scalable estimation methods are available for other models, including stochastic block models (Vu et al., 2013), models which assume that links are independent conditional on latent structure (Mele et al., 2023; Athreya et al., 2018; Kitamura and Laage, 2024), and models for undirected networks (Dahbura et al., 2021; Babkin et al., 2020; Schweinberger and Stewart, 2020), no scalable method for models of directed networks is implemented in publicly available and user-friendly software.
To appreciate the computational challenges involved in estimating directed network formation models, note that the loglikelihood function of the parameter vectors $\bm{\theta}$ and $\bm{\eta}$, given a single observation of a software dependency network $\bm{g}$, is
$\begin{array}[]{lll}\ell(\bm{\theta},\,\bm{\eta})&\coloneqq&\displaystyle\log\sum_{\bm{z}\in\mathcal{Z}}p\left(\bm{z};\,\bm{\eta}\right)\,\pi(\bm{g};\,\bm{x},\,\bm{z},\,\bm{\theta}),\end{array}$ (15)
where $\mathcal{Z}$ is the set of all $K^{N}$ possible configurations of types $\bm{z}$,  $p$ is the multinomial probability of $\bm{z}$ according to (1),  and $\pi$ is the stationary distribution of $\bm{g}$ conditional on observables $\bm{x}$, unobservables $\bm{z}$, and the structural parameter vector $\bm{\theta}$. Maximizing $\ell$ with respect to $\bm{\theta}$ and $\bm{\eta}$ is challenging, because it involves a sum over all $K^{N}$ possible assignments of $N$ packages to $K$ types, which is exponential in $N$. Worse, for each of the $K^{N}$ possible assignments of $N$ packages to $K$ types, the stationary distribution $\pi$ is a function of an intractable normalizing constant, which is the sum of between- and within-type sums, each of which is exponential in the number of pairs of packages $\binom{N_{k}}{2}$ of type $k$ ($k=1,\ldots,K$). Therefore, direct computation of the likelihood function is infeasible unless $N$ is small (e.g., $N\leq 10$).
To develop scalable methods for directed network formation models, we follow Babkin et al. (2020) and decouple the estimation of the unobserved types from the structural utility parameters using a scalable two-step algorithm:
  1. Step 1: Estimate the types of packages.
  2. Step 2: Estimate the structural parameter vector $\bm{\theta}$ conditional on the estimated types.
Step 1
In the first step, we estimate the types of the packages by assuming that there are no externalities ($\gamma=0$) and bounding $\ell$ from below using variational methods, and then using maximization-minorization (MM) algorithm to increase the lower bound on $\ell$. We then assign each package to the type with the highest posterior probability.
We show in Appendix A that the loglikelihood function $\ell$ can be approximated by the following variational lower bound:
$\begin{array}[]{lllll}\ell_{LB}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\,\bm{\Xi})\coloneqq\displaystyle\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{K}\sum_{l=1}^{K}\xi_{ik}\,\xi_{jl}\log\,\pi_{ij,kl}(g_{ij},\bm{x})+\displaystyle\sum_{i=1}^{N}\sum_{k=1}^{K}\xi_{ik}\left(\log\eta_{k}-\log\xi_{ik}\right),\end{array}$ (16)
where $\bm{\Xi}\coloneqq(\xi_{ik})\in[0,1]^{N\times K}$ are auxiliary parameters that serve as approximations of the posterior probabilities that packages $i$ are of type $k$,  $\eta_{k}$ is the probability of being of type $k$,  and $\pi_{ij,kl}$ is the conditional choice probability of a dependency of package $i$ on package $j$, provided $i$ is of type $k$ and $j$ is of type $l$. It is worth noting that $\ell_{LB}$ provides a family of lower bounds on the loglikelihood function $\ell$: Each possible value of the auxiliary parameters $\bm{\Xi}$ provides a lower bound on the loglikelihood function $\ell$. To obtain the best bound on the loglikelihood function $\ell$, we maximize $\ell_{LB}$ with respect to $\bm{\Xi}$.
Maximizing $\ell_{LB}$ with respect to $\bm{\Xi}$ can be cumbersome and time-consuming. Worse, maximization algorithms may be trapped in local maxima of $\ell_{LB}$. Therefore, we construct a surrogate function $M$, which is more convenient to maximize than $\ell_{LB}$:
$M\left(\bm{\Xi};\,\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\Xi}^{(t)}\right)$$\coloneqq$ $\sum_{i=1}^{N}\sum_{j\neq i}^{N}\sum_{k=1}^{K}\sum_{l=1}^{K}\left(\xi_{ik}^{2}\frac{\xi_{jl}^{(t)}}{2\xi_{ik}^{(t)}}+\xi_{jl}^{2}\frac{\xi_{ik}^{(t)}}{2\xi_{jl}^{(t)}}\right)\log\pi_{ij,kl}^{(t)}(g_{ij},\bm{x})$
$+$ $\sum_{i=1}^{N}\sum_{k=1}^{K}\xi_{ik}\left(\log\eta_{k}^{(t)}-\log\xi_{ik}^{(t)}-\frac{\xi_{ik}}{\xi_{ik}^{(t)}}+1\right),$
where $\xi_{ij}^{(t)}$ denotes the estimate of $\xi_{ij}$ at iteration $t$. Relying on the inequality of arithmetic and geometric means and convexity of the negative logarithm, one can show that (4.1) minorizes (16) in the point $\bm{\Xi}^{(t)}$ (see Vu et al., 2013 for further details). Therefore, maximizing (4.1) with respect to $\bm{\Xi}$ forces $\ell_{LB}$ uphill. This step amounts to a quadratic programming problem under box constraints (Stefanov, 2004). This approach along with fast sparse matrix operations, speeds up computations compared with naive variational algorithms that attempt to maximize $\ell_{LB}$ as a function of $\bm{\Xi}$ directly. [^23] See Appendix A for technical details and explanations on how these approximations improve speed of computation. Using these approximations, we can estimate unobserved types at scale while preserving the correlation structure of software dependencies arising from externalities.
Step 2
In the second step, we estimate the structural parameter vector $\bm{\theta}$ by maximizing the pseudolikelihood, which is the product of conditional choice probabilities of links conditional on the types estimated in the first step (Babkin et al., 2020; Martínez Dahbura et al., 2023).
The two-step algorithm sketched above is scalable, in the sense that it scales from small to large networks, and has been tested on undirected networks with a quarter of million units (Dahbura et al., 2021). We provide additional theoretical and computational details in Appendix A and the Online Appendix.
4.2 Results
Table 5: Parameter estimates and standard errors.
Within Between
Estimate Estimate
(Std. Error) (Std. Error)
Edges ($\alpha_{1}$) $-5.801$ $-7.979$
(0.276) (0.006)
$\log(n)\times$ Edges $\left(\alpha_{1}^{\ell}\right)$ $-0.533$
Externality ($\gamma$) $-0.095$
Maturity ($\beta_{1}$) 1.791 $-0.222$
(0.305) (0.011)
$\log(n)\times$ Maturity $\left(\beta_{1}^{\ell}\right)$ $-0.123$
Popularity ($\beta_{2}$) 1.944 0.118
(0.305) (0.01)
$\log(n)\times$Popularity $\left(\beta_{2}^{\ell}\right)$ $-0.18$
Size ($\beta_{3}$) 1.114 $-0.07$
(0.31) (0.011)
$\log(n)\times$Size$\left(\beta_{3}^{\ell}\right)$ $-0.122$
We estimate a model with $K=10$ unobserved types and initialize the types allocation for our algorithm using the Walktrap algorithm. [^24] As suggested by Wainwright and Jordan (2008) and others, we ran the algorithm with multiple initial guesses of the type allocations, but report the results that achieved the highest lower bound. While there is no established method to infer $K$ in this class of models, we have run the estimation with $K=5$ and $K=15$ for robustness, but in both cases the model’s fit was worse than with $K=10$.
The results of our estimation are reported in Table LABEL:tbl:results_python. The first column shows estimates of the structural parameters for links of the same (unobservable) type, while the remaining column provides the estimates for links of different types.
We estimate the externality ($\gamma$) to be negative, providing empirical evidence of the interdependence of links suggested by our strategic model. Essentially, our results provide evidence that when developers form connections, they consider the indirect impacts of their actions, at least partially. This behavior leads to a network architecture that might not be as efficient as one where a central planner aims to maximize the overall utility of all participants. Indeed, the externality encourages developers to establish fewer connections than would be ideal from an aggregate utility-maximization perspective. Maintainers see the number of dependencies of a package they consider linking to as a cost, thus leading to linking decisions that decrease the density of the aggregate network. The negative externality has further implications for contagion, as we explain in the next section.
The parameters $\alpha_{1}$ and $\alpha_{1}^{\ell}$ govern the density of the network and are both estimated negative. The parameter $\alpha_{1}$ governing links between types is likewise estimated negative. The magnitudes of the estimated parameters imply that—other things being equal—in equilibrium sub-networks of libraries of the same type tend to be less sparse than between-types. The sparsity increases with the number of libraries belonging to a particular type. One interpretation is that $\alpha$ measures costs of forming links, which is higher when there are more available libraries to choose from, possibly because of the need to monitor and evaluate more code when forming a decision.
By contrast, the parameters $\beta_{1},\beta_{2},\beta_{3}$ are all positive within types, indicating that developers are more likely to link to packages of the same type of similar maturity, popularity and size. On the other hand, this homophily seems attenuated by the size of the type, as shown by the negative coefficients of the parameters $\beta_{1}^{\ell},\beta_{2}^{\ell},\beta_{3}^{\ell}$.
When forming dependencies with packages of a different type, the effect of the observables is negative for maturity and popularity, while positive for size. We interpret this as evidence of complementarities across types.
One interpretation consistent with these results is that developers use popular features of same-type packages to include in their own code but use infrastructure-style code from different-type packages, which is not necessarily very popular.
4.3 Model Fit
To assess model fit, we follow the methodology developed in the literature on ERGM and HERGMs and simulate the model under the estimated parameters (Hunter et al., 2008; Babkin et al., 2020). We compute network statistics on those simulated networks and compare them with the corresponding statistics evaluated on the observed network. If the model fits the data, most of the simulated network statistics should resemble their observed counterparts.
In Figure 2 we show a boxplot with summaries of the networks obtained from 100 simulations of the model. The summaries show in Figure 2 are natural summaries, because they are the sufficient statistics of the model conditional on the types of packages.  [^25] We start the simulations at the observed network, and we use a Metropolis-Hastings sampler to update the network. We first run a burn-in of 500,000 iterations, and then sample a network every 50,000 iterations. These summaries are normalized, so that a value of zero indicates a perfect match. The red line is the observed network, so a perfect fit corresponds to all the network statistics concentrated around the red line. The boxplots show that the model is able to capture most of the aggregate structural properties of the Python dependency network: most of the simulated statistics are centered around the observed network. The only exception is the externality (i.e. the number of two-paths in our model), however the observed network falls in the $95\%$ confidence interval of the simulation output. We conclude that the estimated model does a good job in reproducing important features of the network.
Figure 2: Model fit: Network statistics
In Figure 3 we report similar boxplots for in-degree and out-degree, showing that the simulated networks replicate the observed one quite closely. These results are particularly striking as it is usually quite hard to fit such a model for a very large network.
Figure 3: Model Fit: Degree distributions
In the Online Appendix C, we compare the fit of our model to alternative approaches. An important question is whether the unobserved heterogeneity included in our model is needed. To answer this question we compare our estimated model to a model without unobserved types ($K=1$). This corresponds to a standard Exponential Random Graph model (ERGM), as studied in Mele (2017). The ERGM is not able to match the aggregate density of the network, while our model’s simulations are concentrated around the observed number of links. The second question is whether the externalities are important in explaining the topology of the equilibrium network. Therefore, we compare our model to a Stochastic blockmodel with covariates and $K=10$ types. This corresponds to our model with $\gamma=0$. Our simulations show that the our model does a better job matching the externality within types, and should therefore be preferred. Finally, while there is no currently established method to choose the optimal number of types for this class of complex models, we have done a significant effort in setting an appropriate $K$ for estimation. Our practical approach consists of setting the number of types by checking the model’s fit for different values of $K$. In appendix C we show the fit with $K=5$ and $K=15$. These alternative models overestimate or underestimate the externality level, so our model with $K=10$ is preferred.
5 Contagious Vulnerabilities
A growing literature explores contagion on networks, including the statistical literature on infectious diseases (e.g., the contagion of HIV and coronaviruses). Such models (see, e.g., Schweinberger et al., 2022; Groendyke et al., 2012; Britton and O’Neill, 2002) usually first use a network formation model to generate a network of contacts among agents and then study how an infectious disease following a stochastic process—like the susceptible-infected-recovered model (Kermack and McKendrick, 1972; Hethcote, 2000)—spreads within this network.
We adapt these ideas to the contagion of vulnerabilities in software dependency graphs. The prevalence of vulnerabilities in software motivated the creation of the Common Vulnerabilities and Exposure (CVE) program and the National Vulnerability Database (NVD) by the National Institute of Standards and Technology, which is an agency of the United States Department of Commerce (NIST, 2024). A software package depending on a vulnerable package is potentially vulnerable itself. To see how, consider, for example, the Equifax data breach (Federal Trade Commission, 2024) in 2017, caused by a vulnerability in the popular open source web application framework Apache Struts 2.262626Apache Struts 2 did not properly validate the Content-Type header of incoming http requests when processing file uploads, which allowed attackers to create http requests that include malicious code. Due to the lack of content validation, this allowed attackers to execute arbitrary code on the web server, ultimately granting them full control of the server. For details see the official NIST announcement on CVE-2017-5638 (https://nvd.nist.gov/vuln/detail/CVE-2017-5638). Equifax did not itself develop or maintain Apache Struts 2, but instead used it as part of its own code base. The vulnerability allowed attackers to execute malicious code providing access to the server hosting the web application developed by Equifax using Apache Struts 2 and costing the company more than \$1.5 Billion to date. [^27] See https://www.bankinfosecurity.com/equifaxs-data-breach-costs-hit-14-billion-a-12473. Article accessed 27 January 2024.
Package managers are part of a large suite of tools that help systems administrators to keep their systems up to date and patch any known vulnerabilities. It takes time for vulnerable systems to be updated (Edgescan, 2022), [^30] It takes organizations between 146 and 292 days to patch cyber security vulnerabilities (Source: missing
6 Conclusion
The network of dependencies among software packages is an interesting laboratory to study network formation and how incentives and externalities shape the topology of the network. Indeed, the creation of modern software gains efficiency by re-using existing libraries; on the other hand, dependencies expose new software packages to bugs and vulnerabilities from other libraries. This feature of the complex network of dependencies motivates the interest in understanding the incentives and equilibrium mechanisms driving the formation of such networks.
In this paper, we estimate a directed network formation model to undertake a structural analysis of the motives, costs, benefits and externalities that a maintainer faces when developing a new software package. The empirical model allows us to disentangle observable from unobservable characteristics that affect the decisions to form dependencies to other libraries.
Using data from the Python dependency network, we find evidence that coders create negative externalities for other coders when creating a link. This raises more questions about the formation of dependency graphs. Using our estimated model, we study the vulnerability of the dependency network to spread of a bug. We measure systemic risk as the average number of downstream packages that are affected by a vulnerability, assuming that the contagion follows a SIR model of epidemic. We simulate interventions that target nodes according to their in-degree, expected fatality and betweenness centrality. Such interventions are unlikely to significantly decrease the risk profile of the network. On the other hand, the increase in AI-assisted tools for programmers, may have a beneficial effect by reducing the cost of producing code in-house, without much need for dependencies.
While we have focused on a stationary realization of the network, there are important dynamic considerations in the creation of these dependencies. While the modeling of forward-looking maintainers may be useful to develop intuition about intertemporal strategic incentives and motives, we leave this development to future work, as it involves several complications in the estimation. Finally, we have focused on a single language, but the analysis can be extended to other languages as well, such as Java, R, Rust, C++ and others.
Appendix A Computational details for estimation
Estimation of our model is challenging because of the normalizing constants $c_{kk}$ in the likelihood and the discrete mixture model for the block assignments. We bypass these issues by extending recently developed approximate two-step estimation methods to directed networks, estimating the structural parameters in two steps.
Formally, the full likelihood of our model can be written as follows
$\mathcal{L}(\bm{g},\bm{x};\bm{\alpha},\bm{\beta},\gamma,\bm{\eta})=\sum_{\bm{z}\in\mathcal{Z}}L\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\gamma,\bm{\eta}\right)=\sum_{\bm{z}\in\mathcal{Z}}p_{\bm{\eta}}\left(\bm{Z}=\bm{z}\right)\,\pi(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\gamma).$ (18)
where each package’s type is i.i.d. multinomial, so the distribution $p_{\bm{\eta}}(\bm{z})$ is
[…]
Since maximizing (18) involves a sum over all possible type allocations of each node as well as a nested normalizing constant in $\pi(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\gamma)$, evaluating the function is infeasible. To ameliorate this issue, we divide estimation into two parts. First, we estimate the type of each node and, second, we estimate the parameters $\bm{\alpha},\bm{\beta},$ and $\gamma$ conditional on the type of each node. This two-step algorithm exploits the fact that our model corresponds to a stochastic blockmodel with covariates when $\gamma=0$, i.e. when there are no link externalities.
A.1 STEP 1: Approximate estimation of unobserved block structure
A.1.1 Variational EM algorithm with MM updates
To recover the types and the block structure of the network, we approximate the model using a stochastic blockmodel. Define
$L_{0}\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\bm{\eta}\right):=p_{\eta}(\bm{z})\pi(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\gamma=0)$ (19)
as the likelihood of a stochastic blockmodel, which means that
  1. 1.
    Each node belongs to one of $K$ blocks/types;
  2. 2.
    Each link is conditionally independent, given the block structure
Then under some conditions – namely, that the network is large, each block/type is not too large with respect to the network, and the employed statistics are in some sense well-behaved (Babkin et al., 2020; Schweinberger, 2020; Schweinberger and Stewart, 2020) – we have
$L\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\gamma,\bm{\eta}\right)\approx L_{0}\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\bm{\eta}\right).$ (20)
Denoting by $q_{\xi}(\bm{z})$ a the auxiliary distribution characterized by parameter $\bm{\Xi}$ approximating the distribution $p_{\eta}(\bm{z})$, the log-likelihood of our model stated in (18) can be lower bounded as follows:
$\ell(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\gamma,\bm{\eta})$ ::=:= $\log\sum_{\bm{z}\in\mathcal{Z}}L\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\gamma,\bm{\eta}\right)$ (21)
$\approx$ $\log\sum_{\bm{z}\in\mathcal{Z}}L_{0}\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\bm{\eta}\right)$
$=$ $\log\sum_{\bm{z}\in\mathcal{Z}}q_{\xi}(\bm{z})\frac{L_{0}\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\bm{\eta}\right)}{q_{\xi}(\bm{z})}$
$\geq$ $\sum_{\bm{z}\in\mathcal{Z}}q_{\xi}(\bm{z})\log\left(\frac{L_{0}\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\bm{\eta}\right)}{q_{\xi}(\bm{z})}\right)$
$=$ $\sum_{\bm{z}\in\mathcal{Z}}q_{\xi}(\bm{z})\log L_{0}\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\bm{\eta}\right)-\sum_{\bm{z}\in\mathcal{Z}}q_{\xi}(\bm{z})\log q_{\xi}(\bm{z})$
$=$ $\mathbb{E}_{q}\log L_{0}\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\bm{\eta}\right)+\mathbb{H}(q)$
$\eqqcolon$ $\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\Xi}).$
where
$\mathbb{H}(q)=-\sum_{\bm{z}\in\mathcal{Z}}q_{\xi}(\bm{z})\log q_{\xi}(\bm{z})$
is the entropy of auxiliary distribution $q_{\xi}(\bm{z})$. In the third row, we multiply and divide by $q_{\xi}(\bm{z})$ and apply Jensen’s inequality in the consecutive line.
The best lower bound is obtained by choosing $q_{\xi}(z)$ from the set of distributions $\mathcal{Q}$ that solves the following variational problem
$\ell(\bm{\alpha},\bm{\beta},\bm{\eta})=\sup_{\bm{\Xi}\,\in\,[0,1]^{N\times K}}\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\Xi})$
However, this variational problem is usually intractable unless we impose more structure on the problem. In practice, researchers restrict the set $\mathcal{Q}$ to a smaller set of tractable distributions (Wainwright and Jordan, 2008; Mele and Zhu, 2023). In the case of the stochastic blockmodel, it is useful and intuitive to restrict $\mathcal{Q}$ to the set of multinomial distributions
[…]
where we collect the variational parameters in the matrix $\bm{\Xi}=(\xi_{ik})\in[0,1]^{N\times K}$. This leads to the following tractable lower bound of $\ell$
$\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\Xi})$$\equiv$ $\sum_{\bm{z}\in\mathcal{Z}}q_{\bm{\Xi}}(\bm{z})\log\left(\frac{L_{0}\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\bm{\eta}\right)}{q_{\bm{\Xi}}(\bm{z})}\right)$
$=$ $\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{K}\sum_{l=1}^{K}\xi_{ik}\xi_{jl}\log\,\pi_{ij,kl}(g_{ij},\bm{x})$
$+$ $\sum_{i=1}^{N}\sum_{k=1}^{K}\xi_{ik}\left(\log\eta_{k}-\log\xi_{ik}\right),$
where the function $\pi_{ij,kl}$ is the conditional probability of a link between $i$ and $j$ of types $k$ and $l$, respectively,
$\log\,\pi_{ij,kl}(g_{ij},\bm{x})$$\coloneqq$ $g_{ij}\log\left(\frac{\exp\left(u_{ij,kl}(\bm{\alpha},\bm{\beta})+u_{ji,lk}(\bm{\alpha},\bm{\beta})\right)}{1+\exp\left(u_{ij,kl}(\bm{\alpha},\bm{\beta})+u_{ji,lk}(\bm{\alpha},\bm{\beta})\right)}\right)$
$+$ $(1-g_{ij})\log\left(\frac{1}{1+\exp\left(u_{ij,kl}(\bm{\alpha},\bm{\beta})+u_{ji,lk}(\bm{\alpha},\bm{\beta})\right)}\right)$
and
$u_{ij,kl}(\bm{\alpha},\bm{\beta})=u(\bm{x}_{i},\bm{x}_{j},z_{ik}=z_{jl}=1,\bm{z};\bm{\alpha},\bm{\beta}).$
Note that $g_{ii}\coloneqq 0$ for all $i$ by definition. The covariate information $\bm{x}$ encodes in our setting categorical information on the level of repositories (maturity, popularity, and size). We, therefore, introduce for the pair of repositories $(i,j)$ the covariate vector $\bm{x}_{ij}=(x_{ij,1},\ldots,x_{ij,p})$, where the $k$th entry is a binary indicator whether repositories $i$ and $j$ match on the $k$th covariate. The variables $x_{p,ij}\coloneqq\bm{1}\{x_{ip}=x_{jp}\}$ are, therefore, indicators for homophily. Generalizations are possible, as long as we maintain the discrete nature of the covariates $\bm{x}$. Including continuous covariates is possible, but this would result in a significant slowdown of this algorithm. Observing (A.1.1), we can drop the dependence on the particular pair of repositories and write $\pi_{kl}(g_{ij},\bm{x}_{ij})=\pi_{ij,kl}(g_{ij},\bm{x})$.
We maximize (A.1.1) in a blockwise manner,first in terms of $\bm{\eta},\bm{\alpha},$ and $\bm{\beta}$ conditional on $\bm{\Xi}$ and then the other way around. Denoting the values of $\bm{\Xi},\bm{\eta},\bm{\alpha},$ and $\bm{\beta}$ in the $t$th iteration by $\bm{\Xi}^{(s)},\bm{\eta}^{(s)},\bm{\alpha}^{(s)},$ and $\bm{\beta}^{(s)}$, the two steps comprising the Variational Expectation Maximization algorithm are:
  • Step 1: Given $\bm{\Xi}^{(s)}$, find $\bm{\alpha}^{(s+1)}$ and $\bm{\beta}^{(s+1)}$ satisfying:
    $\ell_{B}\left(\bm{g},\bm{x},\bm{\alpha}^{(s+1)},\bm{\beta}^{(s+1)},\bm{\eta}^{(s+1)};\bm{\Xi}^{(s)}\right)\,\geq\,\ell_{B}\left(\bm{g},\bm{x},\bm{\alpha}^{(s)},\bm{\beta}^{(s)},\bm{\eta}^{(s)};\bm{\Xi}^{(s)}\right);$
  • Step 2: Given $\bm{\alpha}^{(s+1)}$ and $\bm{\beta}^{(s+1)}$, find $\bm{\Xi}^{(s)+1}$ satisfying:
    $\ell_{B}\left(\bm{g},\bm{x},\bm{\alpha}^{(s+1)},\bm{\beta}^{(s+1)},\bm{\eta}^{(s+1)};\bm{\Xi}^{(s+1)}\right)\,\geq\,\ell_{B}\left(\bm{g},\bm{x},\bm{\alpha}^{(s+1)},\bm{\beta}^{(s+1)},\bm{\eta}^{(s+1)};\bm{\Xi}^{(s)}\right).$
Step 1:
Taking first order conditions with respect to each parameter implies the following closed-form update rules for $\bm{\eta}$ and $\pi_{kl}^{(s+1)}(d,x_{1},\ldots,x_{p})$:
$\eta_{k}^{(s+1)}:=\frac{1}{n}\sum_{i=1}^{n}\xi_{ik}^{(s)},\quad k=1,\ldots,K,$ (25)
and
$\pi_{kl}^{(s+1)}(d,x_{1},\ldots,x_{p}):=\frac{\sum_{i=1}^{n}\sum_{j\neq i}\xi_{ik}^{(s)}\xi_{jl}^{(s)}\bm{1}\{g_{ij}=d,X_{1,ij}=x_{1},\ldots,X_{p,ij}=x_{p}\}}{\sum_{i=1}^{n}\sum_{j\neq i}\xi_{ik}^{(s)}\xi_{jl}^{(s)}\bm{1}\{X_{1,ij}=x_{1},\ldots,X_{p,ij}=x_{p}\}},$ (26)
for $k,l=1,\ldots,K$ and $d,x_{1},\ldots,x_{p}\in\{0,1\}$, respectively. The probability $\pi_{kl}^{(s+1)}(d,x_{1},\ldots,x_{p})$ is the same for all packages $i$ and $j$ such that $z_{i}=k$ and $z_{j}=l$. Therefore, we Note that we are not interested in finding particular estimates for $\bm{\alpha}$, $\bm{\beta}$ and $\gamma$ when running the Variational EM (VEM) algorithm. We are only estimating the probabilities $\pi_{kl}(g,\bm{x})$ for $g\in\{0,1\}$ and $\bm{x}\in\mathcal{X}$, where $\mathcal{X}$ is the set of all possible outcomes of the categorical covariates used in Step 1. This feature allows us to speed up the computation of several orders of magnitude (Dahbura et al., 2021). Finally, remark that updates (25) and (26) can both be comprehended as functions of $\bm{\Xi}^{(s)}$.
Step 2:
For very large networks, maximizing the lower bound in the second step with respect to $\bm{\Xi}$ for given estimates of $\bm{\eta},\bm{\alpha},$ and $\bm{\beta}$ is even under the variational approximation of (22) cumbersome and impractically slow. We thus borrow an idea from Vu et al. (2013) define a function $M\left(\bm{\Xi};\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\Xi}^{(s)}\right)$ that minorizes the lower bound $\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\Xi})$ in $\bm{\Xi}^{(s)}$. Therefore,
$M\left(\bm{\Xi};\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\Xi}^{(s)}\right)$$\leq$ $\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\Xi})\text{ \ \ for all }\bm{\Xi}\in[0,1]^{N\times K}$
$M\left(\bm{\Xi}^{(s)};\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\Xi}^{(s)}\right)$ $=$ $\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\Xi}^{(s)})$
holds with fixed $\bm{\alpha},\bm{\beta},\bm{\eta}$ and $\bm{\Xi}^{(s)}$. Applying these two properties guarantees that maximizing the function $M$ forces the lower bound uphill:
$\begin{array}[]{llllllllll}\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\Xi}^{(s)})&=&M\left(\bm{\Xi}^{(s)};\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\Xi}^{(s)}\right)&\leq&M\left(\bm{\Xi}^{(s+1)};\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\Xi}^{(s)}\right)\\ &\leq&\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\Xi}^{(s+1)})\end{array}$
Extending the approach of Vu et al. (2013) to directed networks, the following expression minorizes $\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\Xi})$
$M\left(\bm{\Xi};\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\Xi}^{(s)}\right)$ ::=:= $\sum_{i=1}^{N}\sum_{j\neq i}^{N}\sum_{k=1}^{K}\sum_{l=1}^{K}\left(\xi_{ik}^{2}\frac{\xi_{jl}^{(s)}}{2\xi_{ik}^{(s)}}+\xi_{jl}^{2}\frac{\xi_{ik}^{(s)}}{2\xi_{jl}^{(s)}}\right)\log\,\pi_{kl}^{(s)}(g_{ij},\bm{x}_{ij})$
$+$ $\sum_{i=1}^{N}\sum_{k=1}^{K}\xi_{ik}\left(\log\eta_{k}^{(s)}-\log\xi_{ik}^{(s)}-\frac{\xi_{ik}}{\xi_{ik}^{(s)}}+1\right).$
In Section B, we detail how sparse matrix multiplication can be exploited to speed up the updates of $\bm{\Xi}^{(s+1)}$.
We run this algorithm until the relative increase of the variational lower bound in below a certain threshold to obtain estimates for $\widehat{\bm{\Xi}}$ and $\widehat{\bm{\eta}}$. We then assign each node to its modal estimated type. Therefore, package $i$ is of type $k$, $\widehat{z}_{ik}=1$, if $\widehat{\bm{\Xi}}_{ik}\geq\widehat{\bm{\Xi}}_{i\ell}$ for all $\ell\neq k$ and for all $i$s.
ONLINE APPENDIX
Appendix B Exploiting sparse matrix operations for updating variational parameters
Next, we detail how the updates of $\mathbf{\xi}$ in Step 1 can be carried out in a scalable manner based on sparse matrix operations. Equation (A.1.1) can be written as a quadratic programming problem in $\mathbf{\xi}$ with the block constraints that $\sum_{k=1}^{K}\xi_{ik}=1$ needs to hold for all $i=1,\ldots,N$. We rearrange the first part of (A.1.1) as follows:
$\sum_{i=1}^{N}\sum_{j\neq i}^{N}\sum_{k=1}^{K}\sum_{l=1}^{K}\left(\xi_{ik}^{2}\,\frac{\xi_{jl}^{(s)}}{2\xi_{ik}^{(s)}}+\xi_{jl}^{2}\,\frac{\xi_{ik}^{(s)}}{2\xi_{jl}^{(s)}}\right)\log\,\pi_{kl}^{(s)}(g_{ij},\bm{x}_{ij})$
$=\sum_{i=1}^{N}\sum_{k=1}^{K}\sum_{j\neq i}^{N}\sum_{l=1}^{K}\xi_{ik}^{2}\,\frac{\xi_{jl}^{(s)}}{2\xi_{ik}^{(s)}}\left(\log\,\pi_{kl}^{(s)}(g_{ij},\bm{x}_{ij})+\log\,\pi_{lk}^{(s)}(g_{ji},\bm{x}_{ji})\right)$
$=\sum_{i=1}^{N}\sum_{k=1}^{K}\frac{\Omega_{ik}(\bm{g},\bm{x},\bm{\Xi}^{(s)})}{2\xi_{ik}^{(s)}}\,\xi_{ik}^{2}$
with
$\Omega_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)~{}\coloneqq~{}\sum_{j\neq i}^{N}\sum_{l=1}^{K}\xi_{jl}^{(s)}\left(\log\,\pi_{kl}^{(s)}(g_{ij},\bm{x}_{ij})+\log\,\pi_{lk}^{(s)}(g_{ji},\bm{x}_{ji})\right).$ (28)
This yields for (A.1.1)
$M\left(\bm{\Xi};\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\Xi}^{(s)}\right)$ $=\sum_{i=1}^{N}\sum_{k=1}^{K}\left(\frac{\Omega_{ik}^{(s)}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)}{2\,\xi_{ik}^{(s)}}-\frac{1}{\xi_{ik}^{(s)}}\right)\xi_{ik}^{2}+\left(\log\eta_{k}^{(s)}-\log\xi_{ik}^{(s)}+1\right)\xi_{ik}$
$=\sum_{i=1}^{N}\sum_{k=1}^{K}A_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)\,\xi_{ik}^{2}+B_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)\,\xi_{ik}$
where
$A_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)\coloneqq\frac{\Omega_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)}{2\,\xi_{ik}^{(s)}}-\frac{1}{\xi_{ik}^{(s)}}$
is the quadratic term and
$B_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)\coloneqq\log\eta_{k}^{(s)}-\log\xi_{ik}^{(s)}+1$
the linear term of the quadratic problem.
To update the estimate of $\bm{\Xi}$, we need to evaluate $A_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ and $B_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ for $i=1,\ldots,N$ and $k=1,\ldots,K$, which, when done naively, is of complexity $O(N^{2}\,K^{2})$. Computing $\Omega_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ for $A_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ is the computational bottleneck driving these complexities. It is thus prohibitively difficult for a large number of population members and communities in the network. Note that in our application, this problem is exasperated as we have more than 35,000 nodes and our theoretical result on the adequacy of the estimation procedure assumes that $K$ grows as a function of $N$. To avoid this issue, we show how $\Omega_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ can be evaluated through a series of sparse matrix multiplications. Thereby, our implementation is fast and light on memory requirements. The underlying idea is that $\Omega_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ collapses to a simple form if $\mathbf{g}$ is an empty network with all pairwise covariates set to zero, i.e., $g_{ij}=0$ and $x_{q}=0\,\forall\,i,j=1,\ldots,N$ and $q=1,\ldots,p$. Given that, in reality, these are seldom the observed values, we, consecutively, go through all connections where either $g_{ij}=1$ or $X_{ij,1}=x_{1},\ldots,X_{ij,p}=x_{p}$ for $\sum_{q=1}^{p}\,x_{q}\neq 0$ holds and correct for the resulting error. In formulae, this implies the following:
$\Omega_{ik}(\bm{g},\bm{x},\bm{\Xi}^{(s)})~{}$ $\coloneqq~{}\sum_{j\neq i}^{N}\sum_{l=1}^{K}\xi_{jl}^{(s)}\left(\log\,\pi_{kl}^{(s)}(g_{ij},\bm{x}_{ij})+\log\,\pi_{lk}^{(s)}(g_{ji},\bm{x}_{ji})\right)$
$=\sum_{j\neq i}^{N}\sum_{l=1}^{K}\xi_{jl}^{(s)}\left(\log\,\pi_{kl}^{(s)}(0,\bm{0})+\log\,\pi_{lk}^{(s)}(0,\bm{0})\right)$
$+\,\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left(g_{ij}\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{kl}^{(s)}(1,\bm{x}_{ij})}{\pi_{kl}^{(s)}(0,\bm{0})}\right)+g_{ji}\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{lk}^{(s)}(1,\bm{x}_{ji})}{\pi_{lk}^{(s)}(0,\bm{0})}\right)\right)$
$+\,\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left((1-g_{ij})\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{kl}^{(s)}(0,\bm{x}_{ij})}{\pi_{kl}^{(s)}(0,\bm{0})}\right)+(1-g_{ji})\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{lk}^{(s)}(0,\bm{x}_{ji})}{\pi_{lk}^{(s)}(0,\bm{0})}\right)\right)$
$=\Omega_{ik}\left(\bm{0},\bm{0},\bm{\Xi}^{(s)}\right)+\Lambda_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right),$ (29)
with
$\Lambda_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ $=\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left(g_{ij}\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{kl}^{(s)}(1,\bm{x}_{ij})}{\pi_{kl}^{(s)}(0,\bm{0})}\right)+g_{ji}\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{lk}^{(s)}(1,\bm{x}_{ji})}{\pi_{lk}^{(s)}(0,\bm{0})}\right)\right)$ (30)
$+\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left((1-g_{ij})\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{kl}^{(s)}(0,\bm{x}_{ij})}{\pi_{kl}^{(s)}(0,\bm{0})}\right)+(1-g_{ji})\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{lk}^{(s)}(0,\bm{x}_{ji})}{\pi_{lk}^{(s)}(0,\bm{0})}\right)\right)$
being the error in $\Omega_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ arising from assuming an empty network with all categorical covariates set to zero. Next, we show how both terms, $\Omega_{ik}\left(\bm{0},\bm{0},\bm{\Xi}^{(s)}\right)$ and $\Lambda_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$, can be computed through matrix operations, completing our scalable algorithmic approach.
We, first, assume an empty network where all categorical covariates set to zero, $g_{ij}=x_{q}=0$ for all $i\neq j$ and $q=1,...,p$ and compute $\Omega_{ik}\left(\bm{0},\bm{0},\bm{\Xi}^{(s)}\right)$ via matrix multiplications. Observe that
$\Omega_{ik}\left(\bm{0},\bm{0},\bm{\Xi}^{(s)}\right)$ $=\sum_{j\neq i}^{N}\sum_{l=1}^{K}\xi_{jl}^{(s)}\left(\log\,\pi_{kl}^{(s)}(0,\bm{0})+\log\,\pi_{lk}^{(s)}(0,\bm{0})\right)$
$=\sum_{l=1}^{K}\sum_{j\neq i}^{N}\xi_{jl}^{(s)}\left(\log\,\pi_{kl}^{(s)}(0,\bm{0})+\log\,\pi_{lk}^{(s)}(0,\bm{0})\right)$
$=\sum_{l=1}^{K}\left(\sum_{j=1}^{N}\xi_{jl}^{(s)}-\xi_{il}^{(s)}\right)\left(\log\,\pi_{kl}^{(s)}(0,\bm{0})+\log\,\pi_{lk}^{(s)}(0,\bm{0})\right)$
$=\sum_{l=1}^{K}\left(\tau^{(s)}_{l}-\xi_{il}^{(s)}\right)\left(\log\,\pi_{kl}^{(s)}(0,\bm{0})+\log\,\pi_{lk}^{(s)}(0,\bm{0})\right)$
holds, where
$\tau^{(s)}_{l}~{}\coloneqq~{}\sum_{j=1}^{N}\xi_{jl}^{(s)}.$
With
$\bm{A_{0}}\left(\bm{\Xi}^{(s)}\right)~{}\coloneqq~{}\begin{bmatrix}\tau^{(s)}_{1}-\xi_{11}^{(s)}&\tau^{(s)}_{2}-\xi_{12}^{(s)}&\dots&\tau^{(s)}_{K}-\xi_{1K}^{(s)}\\ \tau^{(s)}_{1}-\xi_{21}^{(s)}&\tau^{(s)}_{2}-\xi_{22}^{(s)}&\dots&\tau^{(s)}_{K}-\xi_{2K}^{(s)}\\ \vdots&\vdots&\ddots&\vdots\\ \tau^{(s)}(1)-\xi_{n1}^{(s)}&\tau^{(s)}_{2}-\xi_{n2}^{(s)}&\dots&\tau^{(s)}_{K}-\xi_{nK}^{(s)}\end{bmatrix}$
and
$\bm{\Pi_{0}}\left(\bm{\Xi}^{(s)}\right)~{}\coloneqq~{}\begin{bmatrix}\log\,\pi_{11}^{(s)}(0,\bm{0})&\log\,\pi_{12}^{(s)}(0,\bm{0})&\dots&\log\,\pi_{1K}^{(s)}(0,\bm{0})\\ \log\,\pi_{21}^{(s)}(0,\bm{0})&\log\,\pi_{22}^{(s)}(0,\bm{0})&\dots&\log\,\pi_{2K}^{(s)}(0,\bm{0})\\ \vdots&\vdots&\ddots&\vdots\\ \log\,\pi_{K1}^{(s)}(0,\bm{0})&\log\,\pi_{K2}^{(s)}(0,\bm{0})&\dots&\log\,\pi_{KK}^{(s)}(0,\bm{0})\end{bmatrix},$
it follows that
$\bm{A_{0}}\left(\bm{\Xi}^{(s)}\right)\left(\left(\bm{\Pi_{0}}\left(\bm{\Xi}^{(s)}\right)\right)^{\top}+\bm{\Pi_{0}}\left(\bm{\Xi}^{(s)}\right)\right)=\left(\Omega_{ik}\left(\bm{0},\bm{0},\bm{\Xi}^{(s)}\right)\right)_{ik}$ (31)
holds. Put differently, we are able to write $\Omega_{ik}^{(s)}(\bm{0},\bm{0},\bm{\Xi})$ as the $(i,k)$th entry of the result of (31). To compute $\bm{\Pi_{0}}\left(\bm{\Xi}^{(s)}\right)$, we first evaluate the matrix $\pi^{(s)}(1,\bm{0})\in[0,1]^{K\,\times\,K}$ via sparse matrix operations:
$\bm{\pi}^{(s+1)}(1,\bm{0})=\left(\left(\bm{\Xi}^{(s+1)}\right)^{\top}\bm{g}\circ(\bm{J}-\bm{X_{1}})\circ\cdots\circ(\bm{J}-\bm{X_{p}})\ \left(\bm{\Xi}^{(s+1)}\right)\right)\oslash$ (32)
$\left(\left(\bm{\Xi}^{(s+1)}\right)^{\top}(\bm{J}-\bm{X_{1}})\circ\cdots\circ(\bm{J}-\bm{X_{p}})\left(\bm{\Xi}^{(s+1)}\right)\right),$
where $\bm{A}\circ\bm{B}$ denotes the Hadamard (i.e., entry-wise) product of the conformable matrices $\bm{A}$ and $\bm{B}$, $\bm{J}$ is a $N\times N$ matrix whose off-diagonal entries are all one and whose diagonals are all zero. We follow the approach Dahbura et al. (2021) to calculate (32) without breaking the sparsity of the covariate matrices and $\bm{g}$. We can then set
$\bm{\Pi_{0}}\left(\bm{\Xi}^{(s)}\right)=\log\left(1-\bm{\pi}^{(s+1)}(1,\bm{0})\right)$
and calculate $\Omega_{ik}^{(s)}\left(\bm{0},\bm{0},\bm{\Xi}^{(s)}\right)$ for $i=1,\ldots,N$ and $k=1,\ldots,K$.
Next, we correct for the error arising from assuming that $\bm{g}=\bm{0}$ and $\bm{x}=\bm{0}$ holds. Since all covariates are assumed to be categorical, the pairwise covariate vector $\bm{x}_{ij}$ can have at most $2^{p}$ possible outcomes, which we collect in the set $\mathcal{X}$. Given this, one may rewrite the sum over $j\neq i$ as a sum over all possible outcomes of the pairwise covariate information:
$\Lambda_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)~{}$ $=~{}\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left(g_{ij}\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{kl}^{(s)}(1,\bm{x}_{ij})}{\pi_{kl}^{(s)}(0,\bm{0})}\right)+g_{ji}\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{lk}^{(s)}(1,\bm{x}_{ji})}{\pi_{lk}^{(s)}(0,\bm{0})}\right)\right)$
$+\,\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left((1-g_{ij})\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{kl}^{(s)}(0,\bm{x}_{ij})}{\pi_{kl}^{(s)}(0,\bm{0})}\right)+(1-g_{ji})\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{lk}^{(s)}(0,\bm{x}_{ji})}{\pi_{lk}^{(s)}(0,\bm{0})}\right)\right)$
$=\sum_{\bm{x}\in\mathcal{X}}\sum_{j\neq i}^{N}\sum_{l=1}^{K}g_{ij}\,\mathbb{I}(\bm{x}_{ij}=\bm{x})\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{kl}^{(s)}(1,\bm{x})}{\pi_{kl}^{(s)}(0,\bm{0})}\right)$
$+\,g_{ji}\,\mathbb{I}(\bm{x}_{ji}=\bm{x})\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{lk}^{(s)}(1,\bm{x})}{\pi_{lk}^{(s)}(0,\bm{0})}\right)+(1-g_{ij})\,\mathbb{I}(\bm{x}_{ij}=\bm{x})\,\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{kl}^{(s)}(0,\bm{x})}{\pi_{kl}^{(s)}(0,\bm{0})}\right)$
+ $+\,1-g_{ji})\,\mathbb{I}(\bm{x}_{ji}=\bm{x})\,\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{lk}^{(s)}(0,\bm{x})}{\pi_{lk}^{(s)}(0,\bm{0})}\right)$
With
$\bm{\Pi}\left(d,\bm{x},\bm{\Xi}^{(s)}\right)\coloneqq\begin{bmatrix}\log\frac{\pi_{11}^{(s)}(d,\bm{x})}{\pi_{11}^{(s)}(d,\bm{0})}&\log\frac{\pi_{12}^{(s)}(d,\bm{x})}{\pi_{12}^{(s)}(d,\bm{0})}&\dots&\log\frac{\pi_{1K}^{(s)}(d,\bm{x})}{\pi_{1K}^{(s)}(d,\bm{0})}\\ \log\frac{\pi_{21}^{(s)}(d,\bm{x})}{\pi_{21}^{(s)}(d,\bm{0})}&\log\frac{\pi_{22}^{(s)}(d,\bm{x})}{\pi_{22}^{(s)}(d,\bm{0})}&\dots&\log\frac{\pi_{2K}^{(s)}(d,\bm{x})}{\pi_{2K}^{(s)}(d,\bm{0})}\\ \vdots&\vdots&\ddots&\vdots\\ \log\frac{\pi_{K1}^{(s)}(d,\bm{x})}{\pi_{K1}^{(s)}(d,\bm{0})}&\log\frac{\pi_{K2}^{(s)}(d,\bm{x})}{\pi_{K2}^{(s)}(d,\bm{0})}&\dots&\log\frac{\pi_{KK}^{(s)}(d,\bm{x})}{\pi_{KK}^{(s)}(d,\bm{0})}\end{bmatrix}$
and the functions $\bm{\Gamma}\colon\{0,1\}\to\mathbb{R}^{n\times n}$ and $\bm{\Lambda}_{s}\colon\{0,1\}\to\mathbb{R}^{n\times n}$ $(s=1,\ldots,p)$ defined as
$\bm{\Gamma}(d)\coloneqq\begin{cases}\bm{G}&(d=1)\\ \bm{J}-\bm{G}&(d=0)\end{cases}$
and
$\bm{\Delta}(\bm{x})\coloneqq\bm{X_{1}}^{x_{1}}\circ(\bm{J}-\bm{X_{1}})^{1-x_{1}}\circ\cdots\circ\bm{X_{p}}^{x_{p}}\circ(\bm{J}-\bm{X_{p}})^{1-x_{p}},$
we can write
$\Lambda_{ik}$ $\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)=$ $\left(\sum_{\bm{x}\in\mathcal{X}}\sum_{d\in\{0,1\}}\bm{\Gamma}(d)\circ\bm{\Delta}(\bm{x})\,\bm{\Xi}^{(s)}\,\bm{\Pi}^{(s)}\left(d,\bm{x},\bm{\Xi}^{(s)}\right)+\left(\bm{\Gamma}(d)\circ\bm{\Delta}(\bm{x})\right)^{\top}\,\bm{\Xi}^{(s)}\,\bm{\Pi}^{(s)}\left(d,\bm{x},\bm{\Xi}^{(s)}\right)^{\top}\right)_{ik}.$
Extending (32) to generic covariate values $\bm{x}$ yields
$\bm{\pi}^{(s+1)}(1,\bm{x})=$ $\big{(}\left(\bm{\Xi}^{(s+1)}\right)^{\top}\bm{g}\circ\bm{X_{1}}^{x_{1}}\circ(\bm{J}-\bm{X_{1}})^{1-x_{1}}\circ\cdots\circ\bm{X_{p}}^{x_{p}}\circ(\bm{J}-\bm{X_{p}})^{1-x_{p}}\left(\bm{\Xi}^{(s+1)}\right)\big{)}\oslash$
$\left(\left(\bm{\Xi}^{(s+1)}\right)^{\top}\bm{X_{1}}^{x_{1}}\circ(\bm{J}-\bm{X_{1}})^{1-x_{1}}\circ\cdots\circ\bm{X_{p}}^{x_{p}}\circ(\bm{J}-\bm{X_{p}})^{1-x_{p}}\left(\bm{\Xi}^{(s+1)}\right)\right),$
enabling the evaluation of $\bm{\Pi}\left(d,\bm{x},\bm{\Xi}^{(s)}\right)$ for $\bm{x}\in\mathcal{X}$. Following Dahbura et al. (2021), we can still evaluate this matrix by sparse matrix operations.
In sum, we have shown how to compute the terms $\Omega_{ik}\left(\bm{0},\bm{0},\bm{\Xi}^{(s)}\right)$ and $\Lambda_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ through matrix multiplications. Plugging in these terms into (B) enables a fast computation of $\Omega_{ik}(\bm{g},\bm{x},\bm{\Xi}^{(s)})$, which is the computational bottleneck for evaluating the quadratic term needed to update $\bm{\Xi}^{(s)}$ to $\bm{\Xi}^{(s+1)}$.
B.1 STEP 2: Estimation of structural utility parameters
In the second step, we condition on the approximate block structure estimated in the first step $\widehat{\bm{z}}$ and estimate the structural payoff parameters. We compute the conditional probability of a link within types and between types
$p_{ij}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\gamma;\widehat{\bm{z}})$ $=$ $\begin{cases}\Lambda\left(u_{ij}(\alpha_{w},\bm{\beta}_{w})+u_{ji}(\alpha_{w},\bm{\beta}_{w})+4\gamma\sum_{r\neq i,j}I_{ijr}g_{jr}g_{ir}\right)&\text{ if }\widehat{\bm{z}}_{i}=\widehat{\bm{z}}_{j}\\ \Lambda\left(u_{ij}(\bm{\alpha},\bm{\beta})+u_{ji}(\bm{\alpha},\bm{\beta})\right)&\text{ otherwise }\end{cases}$
where $I_{ijr}=1$ if $\widehat{\bm{z}}_{i}=\widehat{\bm{z}}_{j}=\widehat{\bm{z}}_{r}$ and $I_{ijr}=0$ otherwise; and $\Lambda(u)=e^{u}/(1+e^{u})$ is the logistic function.
The maximum pseudolikelihood estimator (MPLE) solves the following maximization problem
$(\widehat{\bm{\alpha}},\widehat{\bm{\beta}},\widehat{\gamma})$ $=$ $\arg\max_{\bm{\alpha},\bm{\beta},\gamma}\ell_{PL}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\gamma;\widehat{\bm{z}})$
$=$ $\arg\max_{\bm{\alpha},\bm{\beta},\gamma}\;\sum_{i\neq j}^{n}\,\left(g_{ij}\log p_{ij}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\gamma;\widehat{\bm{z}})+(1-g_{ij})\log(1-p_{ij}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\gamma;\widehat{\bm{z}}))\right)$
In practice the estimator maximizes the log of the product of conditional probabilities of linking. The asymptotic theory for this estimator is in Boucher and Mourifie (2017) and Stewart and Schweinberger (2023). It can be shown that the estimator is consistent and asymptotically normal. As long as the estimator for $\bm{z}$ provides consistent estimates, the estimator for the structural parameters is well-behaved.
Appendix C Additional figures and estimates
C.1 Comparison with other models
In Figure A1 we show the simulated statistics generated by a model with $K=1$. In aggregate, this model performs very poorly. This suggests that the unobserved heterogeneity can help fit the model better.
Figure A1: Model fit, ERGM ($K=1$)
We compare our model with $K=10$ unobserved types to a stochastic blockmodel with $K=10$ unobserved types. In essence this is the same as our model, but we exclude the externality. As expected the stochastic blockmodel does a worse job predicting the externality within blocks, as shown in Figure A2.
Figure A2: Model fit, stochastic blockmodel with $K=10$
We also did many robustness checks to choose the number of types $K$. In Figures A3 and A4 we report the boxplot for the fit of models with $K=5$ and $K=15$, respectively.
Figure A3: Model fit, model with $K=5$
Figure A4: Model fit, model with $K=15$
From the graph we can conclude that increasing or decreasing the number of types does not seem to help in estimating the externality.
Table A1: Parameter estimates and standard errors.
Within Between
Estimate Estimate
(Std. Error) (Std. Error)
Edges ($\alpha_{1}$) $-5.703$ $-7.979$
(0.274) (0.006)
$\log(n)\times$ Edges ($\alpha_{2})$ $-0.549$
Maturity ($\beta_{1}$) 1.767 $-0.222$
(0.303) (0.011)
$\log(n)\times$ Maturity ($\beta_{1}$) $-0.12$
Popularity ($\beta_{1}$) 1.071 $-0.07$
(0.308) (0.011)
$\log(n)\times$Popularity ($\beta_{1}$) $-0.117$
Size ($\beta_{1}$) 1.92 0.118
(0.303) (0.01)
$\log(n)\times$Size ($\beta_{1}$) $-0.178$
Table A2: Parameter estimates and standard errors ($K=5$).
Within Between
Estimate Estimate
(Std. Error) (Std. Error)
Edges ($\alpha_{1}$) $-4.893$ $-7.45$
(0.263) (0.006)
$\log(n)\times$ Edges $\left(\alpha_{1}^{\ell}\right)$ $-0.569$
Externality ($\gamma$) $-0.258$
Maturity ($\beta_{1}$) 0.711 $-0.267$
(0.302) (0.012)
$\log(n)\times$ Maturity $\left(\beta_{1}^{\ell}\right)$ $-0.012$
Popularity ($\beta_{1}$) 1.316 0.104
(0.31) (0.011)
$\log(n)\times$Popularity ($\beta_{1}$) $-0.111$
Size ($\beta_{1}$) 1.324 $-0.073$
(0.307) (0.011)
$\log(n)\times$Size $\left(\beta_{1}^{\ell}\right)$ $-0.142$
Table A3: Parameter estimates and standard errors ($K=15$).
Within Between
Estimate Estimate
(Std. Error) (Std. Error)
Edges ($\alpha_{1}$) $-4.181$ $-8.28$
(0.305) (0.006)
$\log(n)\times$ Edges $\left(\alpha_{1}^{\ell}\right)$ $-0.765$
Externality ($\gamma$) 0.135
Maturity ($\beta_{1}$) 3.201 $-0.198$
(0.321) (0.011)
$\log(n)\times$ Maturity $\left(\beta_{1}^{\ell}\right)$ $-0.307$
Popularity ($\beta_{1}$) 1.58 0.117
(0.287) (0.01)
$\log(n)\times$Popularity $\left(\beta_{1}^{\ell}\right)$ $-0.15$
Size ($\beta_{1}$) 0.369 $-0.074$
(0.29) (0.01)
$\log(n)\times$Size $\left(\beta_{1}^{\ell}\right)$ $-0.048$
ERGM Model
Estimate
(Std. Error))
Edges $-7.961$
$(0.007)$
Externality $-0.178$
$(0.002)$
Maturity $-0.131$
$(0.010)$
Size $0.127$
$(0.010)$
Popularity $-0.066$
$(0.010)$
Table A4: Parameter estimates and standard errors (K=1)